### v1

# Is Translation All You Need? A Study on Solving Multilingual Tasks with Large Language Models

Abstract
Large language models (LLMs) have demonstrated strong multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora. Existing works leverage this phenomenon to improve their multilingual performances on NLP tasks. In this work, we extend the evaluation from NLP tasks to real user queries. We find that even though translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios. For culture-related tasks that need deep language understanding, prompting in the native language proves to be more promising since it can capture the nuances related to culture and language. Therefore, we advocate for more efforts towards the development of strong multilingual LLMs instead of just English-centric LLMs.
Is Translation All You Need?
A Study on Solving Multilingual Tasks with Large Language Models
[…]
1 Introduction
Large language models (LLMs) frequently demonstrate the capability to understand and generate text across multiple languages, a skill attributed to their training on vast corpora composed of texts from various languages OpenAI (2023); Shi et al. (2022); Muennighoff et al. (2023); Jiang et al. (2023); Nguyen et al. (2023). However, these datasets are often disproportionately dominated by English content Brown et al. (2020); Chowdhery et al. (2022); Workshop et al. (2023), resulting in an English-centric bias in LLMs. This imbalance can subsequently hinder the models’ proficiency in other languages, often leading to suboptimal performance in non-English contexts Ahuja et al. (2023); Lai et al. (2023); Zhang et al. (2023b).
Figure 1: Illustration of two kinds of LLMs on different tasks. "English-centric LLMs" refers to LLMs trained mainly in English corpora. "Multilingual LLMs" refers to ideal LLMs that are equally capable in all languages.
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020). Recent investigations have expanded this concept by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then proceed to solve problems step by step, leads to better performance.
Figure 2: Examples illustrating how translation can both improve (top) and degrade (bottom) the performance of LLMs. The example in Chinese is drawn from MGSM and the example in Swahili is drawn from M3Exam, respectively. Translation is beneficial when the questions are language-independent. However, for questions requiring deep cultural knowledge, translation can hinder the ability to answer correctly.
Despite these advancements, the methodologies employed across different studies vary considerably, and the specific impact of translation in improving multilingual task performance has not been rigorously analyzed. Furthermore, the focus of these studies tends to be on specific NLP tasks and English-centric LLMs, with less attention given to real-world scenarios of multilingual user requests. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across a spectrum of multilingual contexts.
In this study, we conduct a thorough investigation into the utility of translation with large language models for a variety of scenarios, including multilingual NLP tasks and real-world multilingual user queries. We compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages against several baselines, such as using native language prompts and cross-lingual prompts Huang et al. (2023). Our experiments also span across a diverse set of LLMs and translation tools to ensure the broad applicability of our findings.
Our results demonstrate that simply translating queries into English can already achieve the best results in multiple task categories. Subsequently, we investigate the utility of translation in real-world scenarios by conducting tests with requests from ShareGPT [^1] https://sharegpt.com/. We discover that translation tends to yield better responses for low-resource languages. However, when working with high-resource languages and advanced LLMs, employing prompts in native languages appears to be the more effective strategy. To further analyze such behaviors, we extend our evaluations to LLMs optimized for languages other than English, focusing specifically on culture-related tasks. We observe that for these models, prompts in the native languages yield superior results by capturing the nuances related to culture and language, as illustrated in Figure 2 (b). In summary, as illustrated in Figure 1, while translation can improve the performance of English-centric LLMs, it may not be optimal for all scenarios. In this work, we highlight the need for robust LLMs that are equally proficient across all languages to effectively address a broader range of cases. Therefore, we advocate for more efforts towards the development of multilingual LLMs.
2 Translation for NLP Tasks
In this section, we explore various prompting strategies across multiple languages and LLMs, encompassing a wide spectrum of NLP tasks. This allows us to understand how different prompting methods impact the performance of tasks that theoretically do not depend on any specific language.
2.1 Experiment Setup
Original Question
制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
Native-Basic
{Original Question }
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-Basic
{Original Question }
You should format your final answer as "Answer: <Arabic numeral>".
Native-CoT
{Original Question }
让我们一步步思考。
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-CoT
{Original Question }
Let’s think step by step in English.
You should format your final answer as "Answer: <Arabic numeral>".
XLT
I want you to act as an arithmetic reasoning expert for Chinese.
Request: {Original Question}
You should retell the request in English.
You should do step-by-step answer to obtain a number answer.
You should step-by-step answer the request.
You should tell me the answer in this format ’Answer :’.
Trans-Google
Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Trans-NLLB
To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Table 1: An example of zero-shot prompts for a Chinese problem. For Native-Basic, EN-Basic, Native-CoT, EN-CoT and XLT, we provide the original Chinese question as input and expect an answer in the corresponding format; for Trans-Google and Trans-NLLB, we input the translated question in English, and expect a step-by-step solution in English. To obtain the desirable output format, we instruct the models to output in specific format.
We perform assessments on six benchmarks on reasoning, understanding, and generation tasks that can encapsulate various abilities of LLMs. We follow the task and language selection from Huang et al. (2023) to consider a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu). We categorize languages larger than 1% frequency in Common Crawl [^2] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages as high-resource languages (i.e., de, ru, fr, zh, es, ja, it and vi), and the rest as low-resource languages. We exclude English since we want to evaluate the efficient prompting strategy for non-English tasks. As our prompting strategies involve methods with translators, we do not test on the FLORES-200 dataset.
2.1.1 Tasks
We evaluate the models and prompting strategies on the following tasks: 1) the MGSM benchmark Shi et al. (2022), which tests models on grade school math problems. 2) the XCOPA benchmark Ponti et al. (2020), which presents a premise with two choices to determine causality or effect; 3) the XNLI benchmark Conneau et al. (2018), which assesses Natural Language Inference by judging the relationship between a premise and a hypothesis; 4) the PAWS-X benchmark (Yang et al., 2019) involves identifying paraphrases in two sentences; 5) the MKQA dataset (Longpre et al., 2021), which focuses on open-domain questions; 6) the XL-Sum benchmark Hasan et al. (2021), which involves summarizing news articles. We choose a subset of 9 languages for MKQA and 5 languages for XL-Sum. For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks. More details of the benchmarks can be found in A.1.
For each task, we sample up to 500 examples from the test set for each language. For MGSM, since there are only 250 examples for each language, the entire test set will be used. For generation tasks like MKQA and XL-Sum, we will translate the answer back to the original language if the prompting strategy involves a translator.
2.1.2 Models
We conduct experiments with six LLMs, consisting of one closed-source language model and five open-source language models:
  • ChatGPT [^3] https://platform.openai.com/docs/models/gpt-3-5, which is the most capable and cost-effective model in the GPT-3.5 family and it has been optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
  • Mistral-7B-Instruct (v0.2). This model is the instructed version of Mistral-7B Jiang et al. (2023).
  • Llama-2-13B-chat and Llama-2-70B-chat, which are the chat models in Llama-2 family Touvron et al. (2023). Due to computational resource limitations, we use the AWQ Lin et al. (2023) version for Llama-2-70B-chat.
  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including multilingual tasks Muennighoff et al. (2023).
2.1.3 Prompting Strategies
Considering the language of instruction, whether to apply a chain-of-thought reasoning approach and the availability of translation tools, we assess the following experimental strategies in the experiment. We utilize a zero-shot approach across all strategies, given that the models we selected have been fine-tuned to follow instructions.
Basic prompt with native instructions (Native-Basic)
This is to directly pose the questions without employing additional prompting strategies such as chain-of-thought. Both the query information and the instruction provided are presented in the original language.
Basic prompt with English instructions (EN-Basic)
Compared with Native-Basic, EN-Basic  instructs LLMs with English but the query information is in the original language.
Native chain-of-thought (Native-CoT)
In Native-CoT, we ask the question in the native language and ask the model to reason with the native language with the instruction "Let’s think step by step." translated into that language.
English chain-of-thought (EN-CoT)
We pose the question in the native language but instruct the model to reason in English with the instruction "Let’s think step by step in English".
Cross-lingual-thought (XLT)
XLT Huang et al. (2023) is a state-of-the-art prompting method to handle multilingual NLP tasks. It prompts LLMs to translate the question into English and solve the problem step-by-step in English, leveraging the English-centric bias of LLMs.
Translate to English with Google Translate (Trans-Google)
It uses Google Translate API to translate the original questions into English and then solve the problem step by step.
Translate to English with NLLB models (Trans-NLLB)
Instead of using commercial translators, we use an open-source model, namely NLLB Team et al. (2022). Specifically, we chose nllb-200-3.3B to do the translation.
In addition to the prompting strategies, an output constraint is also included in the template to facilitate answer extraction. The settings are demonstrated in Table 1. In the case of open-source models, where the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
2.2 Results and Discussions
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
ChatGPT Native-Basic 34.4 72.3 52.2 49.7 35.4 31.9 46.0
EN-Basic 41.1 76.1 63.0 62.0 36.5 32.7 51.9
Native-CoT 49.9 72.4 50.6 50.3 35.4 31.1 48.3
EN-CoT 61.1 78.6 56.4 61.6 42.9 29.7 55.1
XLT 62.3 79.3 59.2 59.4 37.6 23.4 53.5
Trans-Google 73.9 84.5 60.5 67.2 43.8 30.3 60.0
Trans-NLLB 61.0 79.7 59.2 67.5 37.2 28.3 55.5
bloomz-7b1 Native-Basic 1.3 21.4 8.3 - 7.8 - 9.7
EN-Basic 2.0 56.5 43.9 - 10.6 - 28.3
Native-CoT 1.2 20.9 8.2 - 6.5 - 9.2
EN-CoT 1.7 53.9 35.9 - 9.3 - 25.2
XLT 1.7 50.5 35.4 - 8.0 - 23.9
Trans-Google 2.7 63.7 44.3 - 17.2 - 32.0
Trans-NLLB 2.4 61.8 43.8 - 14.7 - 30.7
Mistral-7B-Instruct Native-Basic 11.2 54.2 42.8 44.6 7.8 25.9 31.1
EN-Basic 23.8 34.9 50.2 46.9 7.0 21.6 30.7
Native-CoT 17.0 53.8 43.4 44.3 7.8 24.6 31.8
EN-CoT 27.6 40.8 50.0 46.6 11.5 18.7 32.5
XLT 31.8 61.5 46.0 47.8 9.6 17.6 35.7
Trans-Google 41.3 59.2 55.0 51.5 17.0 32.8 42.8
Trans-NLLB 31.7 54.4 53.0 52.4 15.5 32.2 39.9
Llama-2-13b-Chat Native-Basic 15.6 50.1 36.4 54.0 18.3 27.5 33.6
EN-Basic 19.0 54.3 43.4 59.1 20.2 31.1 37.9
Native-CoT 18.1 50.9 35.7 54.8 15.7 27.2 33.7
EN-CoT 19.9 54.5 43.7 57.6 19.8 30.2 37.6
XLT 22.3 56.0 51.4 55.7 19.0 21.2 37.6
Trans-Google 39.1 71.9 46.1 58.4 33.8 35.7 47.5
Trans-NLLB 31.8 68.2 45.4 57.8 28.4 34.2 44.3
Llama-2-70B-Chat Native-Basic 23.6 51.5 39.0 52.8 24.8 29.4 36.2
EN-Basic 28.6 55.3 46.5 60.4 24.7 31.2 40.1
Native-CoT 23.5 50.5 37.9 54.9 21.9 25.8 34.5
EN-CoT 30.2 61.2 45.9 64.9 31.1 30.2 42.3
XLT 32.8 58.7 52.2 55.7 26.6 17.7 38.4
Trans-Google 53.3 80.9 54.0 68.5 39.7 33.8 51.8
Trans-NLLB 43.8 77.1 52.2 69.2 19.4 33.1 49.1
Table 2: Average scores across different languages (excluding English) for the six benchmarks in zero-shot setting. The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks. The best result for each model is in bold.
The main results are shown in Table 2. We also calculate the average performance for high-resource and low-resource languages separately for each task in Table 4 in Appendix A.2. We notice that Trans-Google, despite simple, surprisingly demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Trans-NLLB ranks second on average, trailing behind Trans-Google. The discrepancy between the two methods lies solely in the quality of translation, indicating that the NLLB model’s translation proficiency is inferior. Nonetheless, its open-source nature renders it a potentially favourable choice under specific circumstances. 2) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 3) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. More results and discussions are in Appendix A.2.
3 Translation for Real User Queries
The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages. However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages.
To assess the impact of translation on real-world queries, we extract user requests from ShareGPT datasets. We select 10 languages ranging from high to low resources and randomly sample 50 requests for each language from the first round of the dataset. To evaluate the quality of the responses, we use GPT-4-Turbo (gpt-4-turbo-0125) as the judge. The prompt for the judge is shown in Figure 6 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10. Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we will translate the output back to the original language for consistency. We select the two best models on NLP tasks for evaluation: ChatGPT and Llama-2-70B-Chat.
3.1 Results and Discussions
In our comparison, we analyzed the scores of two sets of responses generated by the same model, calculating the win rate for each language. The results are shown in Figure 3 and we make the following observations: 1) ChatGPT shows different behaviours for high-resource languages and low-resource languages. For Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate. For lower-resource languages, translation is often a better option. 2) For Llama-2-70B-Chat, translation has a higher win rate for all the languages. This echoes the fact that Llama-2-70B-Chat is an English-centric LLM. Even though translation may lose some information, the gain of better understanding after translation dominates the performance.
The distinct behaviours exhibited by Llama-2-70B-Chat and ChatGPT underscore the inherent differences between the LLMs. Llama-2-70B-Chat is predominantly English-centric, as evidenced by its improved performance in both NLP tasks and real user queries when input is translated into English. Conversely, ChatGPT demonstrates characteristics of a “multilingual LLM”, as shown in Figure 1(b), showcasing its broader linguistic capacity. Nevertheless, ChatGPT still relies on translation to bolster its performance for lower-resource languages, indicating room for improvement in its approach to truly multilingual processing.
To ascertain the extent to which answering user queries necessitates an understanding of local cultural nuances, we employed GPT-4-Turbo, utilizing a specially crafted prompt to analyze user queries across multiple languages, as shown in Figure 7 in the Appendix. The results, illustrated in Figure 4, reveal that a significant portion of the queries—ranging from 18% to 50% for each language—require knowledge of local culture for accurate responses. This variation underscores the rich cultural elements embedded within the data, affirming our hypothesis that queries containing culture-specific information pose a greater challenge for the LLMs.
(a) Win rate with ChatGPT
(b) Win rate with Llama-2-70B-Chat
Figure 3: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat.
3.2 Analysis
Figure 4: The percentage of the questions that necessitate local cultural knowledge.
ChatGPT has shown promising performance in some high-resource languages. We extend the experiment to investigate the effectiveness of multilingual LLMs in handling culture-specific tasks, aiming to assess how well these models grasp cultural nuances and idiomatic language. Through our experiment, we seek insights into the adaptability of LLMs across diverse cultural contexts.
3.2.1 Experiment Setup
We select the following models to represent various categories of LLMs and assess the potential differences in their response to translation:
  • ChatGPT and llama-2-70B-chat, which are the strongest models in our previous experiments.
  • Qwen1.5-72B-Chat Bai et al. (2023) and Yi-34B-Chat [^4] https://github.com/01-ai/Yi. These two open-source models demonstrate strong English and Chinese capabilities. We aim to test if they excel more in Chinese than in translating Chinese into English.
Please note that except ChatGPT, the results for the other three models are obtained by calling API from Together [^5] https://api.together.xyz/. We chose open-source models which are larger than 30B since we hypothesise that larger models contain more knowledge.
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions require strong multilingual proficiency and cultural knowledge to perform well. For the subject categories, we choose language and social science, which are more likely to contain language-dependent knowledge. For each language, we select up to 500 samples for evaluation. For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
3.2.2 Results and Discussions
Figure 5: Results of four LLMs on M3Exam (a) language and (b) social science subject categories. Note that in M3Exam, not all subjects are available in every language, resulting in a discrepancy in the number of languages for the two subjects.
Based on the results shown in Figure 5, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance for all languages. This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation process hurts Chinese performance, highlighting the significant influence of translation errors on comprehension. Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat.
From this experiment, it is evident that while translation serves as a straightforward and initially effective strategy to enhance the performance of English-centric LLMs on non-English tasks, it falls short of being optimal. Directly training LLMs in specific languages can yield superior results by eliminating the potential for translation errors.
4 General Discussions
In this work, we categorize multilingual tasks into two groups: language-independent tasks and language-dependent tasks. Among the NLP benchmarks discussed in Section 2, such as MGSM, XCOPA, XNLI, PAWS-X, and MKQA, are parallel benchmarks. This means that the examples maintain semantic equivalence across all languages, categorizing them as language-independent tasks. Language-independent tasks are typically easier to evaluate across different linguistic groups, as they rely on universal cognitive abilities, like reasoning and comprehension, rather than specific linguistic knowledge. In contrast, tasks involving real user queries and culture-related questions, discussed in Section 3, necessitate models to possess cultural and contextual understanding, categorizing them as language-dependent tasks. By distinguishing between language-dependent and language-independent tasks, we aim to promote fair and accurate evaluations across diverse linguistic contexts, enhancing the development and assessment of NLP systems in multilingual environments.
Corresponding to the two types of tasks, multilingual LLMs must develop dual competencies:
General Language Skills:
Essential for understanding and engaging with both task types, these skills enable LLMs to process and comprehend content across multiple languages, facilitating basic instruction interpretation and response generation.
Specific Language Skills:
These skills are crucial for tasks demanding an understanding of cultural nuances and linguistic subtleties, such as interpreting culturally specific queries.
By equipping LLMs with these comprehensive skills, we ensure their effectiveness in a wide array of multilingual contexts, enhancing both linguistic adaptability and cultural sensitivity.
The benefits of translation on NLP tasks demonstrate the dominance of English-centric LLMs. Even top models like ChatGPT, perform better through translation. In handling user queries, ChatGPT excels in native languages compared to translation, particularly in high-resource languages but struggles with low-resource ones. When testing on the M3Exam questions, Qwen1.5-72B-Chat and Yi-34B-Chat perform quite well in the native language for Chinese questions but have no such advantages for other languages. This dependency on translation highlights the challenges in developing LLMs that can equally handle a wide range of languages, especially those with limited resources. Such findings prompt further research and development efforts aimed at enhancing the multilingual efficacy of LLMs, striving for models that can understand and generate text across languages with minimal reliance on translation.
5 Conclusion
In this work, we perform a comprehensive assessment of LLMs across diverse multilingual tasks. These tasks include traditional NLP challenges, real user queries, and culture-specific tasks. We explore the use of translation-based methods as a strategy to overcome the limitations inherent in English-centric LLMs, thereby facilitating enhanced comprehension. However, we find that for tasks heavily reliant on language-specific nuances, employing prompts in the native language emerges as a more effective approach. This method is particularly adept at capturing the subtleties and intricacies unique to each language. Our findings suggest that the development of robust multilingual LLMs presents a more effective solution for addressing the challenges of multilingual tasks, especially for language-dependent tasks. Therefore, we advocate for the prioritization of research and development efforts towards the creation of strong multilingual large language models.
6 Limitations
This study aims to systematically assess the effectiveness of various prompting strategies across different tasks and LLMs. Due to limitations in computing resources, it was not possible to evaluate all existing prompting strategies comprehensively. However, we endeavoured to cover the most commonly employed strategies to formulate a broad conclusion. In our evaluation of LLMs on culture-related tasks, we specifically selected two LLMs optimized for Chinese, acknowledging it as one of the most widely spoken languages globally. The dataset used, M3Exam, comprises exclusively multiple-choice questions. It is important to note this specificity as it may influence the applicability of our findings. In our evaluation, we limited our sampling to up to 500 samples for each language within the benchmarks to manage computational constraints and ensure a broad yet feasible analysis scope. Consequently, our results might not be directly comparable with other studies that evaluate performance across the entire benchmark. In future work, we plan to extend our evaluation to LLMs optimized for other languages and to explore benchmarks presented in various formats beyond multiple-choice questions.
Acknowledgements
This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore. Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and School of Computer Science and Engineering, Nanyang Technological University, Singapore, for their support.
Appendix A Additional Experiments
A.1 NLP Tasks
Here are the detailed descriptions of the NLP tasks:
Arithmetic Reasoning
The MGSM Shi et al. (2022) benchmark includes mathematical problems from grade school and requires the model to compute the accurate solution. It spans 10 languages, and we use the accuracy score for assessment.
Commonsense Reasoning
The XCOPA benchmark Ponti et al. (2020) consists of a single premise and two choices. The goal is to identify which choice is the cause or effect of the premise. It covers 11 languages from various families, with an accuracy score used for evaluation.
Natural Language Inference
The XNLI Conneau et al. (2018) benchmark includes one premise and one hypothesis. The model’s job is to determine if the hypothesis is entailed, contradicted, or neutral based on the premise. It covers 15 languages, and we evaluate it using the accuracy score.
Paraphrase Identification
The PAWS-X (Yang et al., 2019) benchmark consists of two sentences and requires the model to judge whether they are paraphrases. It covers 7 languages, and we assess based on accuracy score.
Question Answering
The MKQA dataset (Longpre et al., 2021) contains open-domain questions that require predicting short answers. Questions that are unanswerable or excessively long to have a specific answer are not considered during evaluation. This dataset covers 25 languages, with our focus on 9 languages: de, es, fr, ja, ru, th, tr, vi, and zh. We assess the model’s performance using the token overlap F1 score.
Summarization
The XL-Sum Hasan et al. (2021) benchmark requires the model to condense a lengthy news article into a brief summary. It covers 44 languages, and we select a subset of 5 languages: es, fr, tr, vi, and zh. We use the ROUGE-1 score for evaluation.
The prompts of EN-Basic for each task are shown in Table 3, which are adapted from Huang et al., 2023. The prompt templates for other prompting strategies and the instructions for output formats are designed according to the descriptions in Section 2.1.3.
A.2 Additional Results
The average performances for high-resource and low-resource languages are shown in Table 4. Table 5, Table 6, Table 7, Table 8, Table 9 and Table 10 shows the detailed results for MGSM, XCOPA, XNLI, PAWS-X, MKQA and XL-Sum, respectively. In addition to the finding in Section 2.2, We find XLT exhibits competitive performance in reasoning tasks; however, its performance in generation tasks is less impressive. Our findings indicate that when employing the XLT prompting strategy, ChatGPT declined to answer 26.4% of the questions in the XL-Sum tasks, responding with “I’m sorry, I cannot …” This refusal pattern was not observed when utilizing other prompting strategies. For open-source models, while we did not observe a refusal pattern, they do not follow the instructions properly, which also degrades their performance with XLT.
Benchmark #Test Basic Prompt
MGSM 250 {problem}
XCOPA 500
Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
XNLI 500
{premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
PAWS-X 500
Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
MKQA 500
Answer the question in one or a few words in {target_language}: {question}?
XL-Sum 500
{article} Summarize the article.
Table 3: Template of EN-Basic for each benchmark. #Test denotes the number of samples in the test set.
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
high low high low high low high low high low high low high low
ChatGPT Native-Basic 44.4 19.4 84.6 69.7 56.9 48.6 51.6 40.6 35.1 36.4 32.5 29.9 50.8 40.8
EN-Basic 50.3 27.3 88.3 73.3 64.6 61.8 64.3 50.4 37.4 33.3 33.3 30.0 56.4 46.0
Native-CoT 65.1 27.1 84.1 69.8 54.9 47.4 51.6 43.4 35.5 35.1 31.9 27.9 53.8 41.8
EN-CoT 70.5 47.1 89.9 75.9 60.2 53.6 63.7 51.2 43.3 41.2 30.0 28.6 59.6 49.6
XLT 70.4 50.1 89.3 76.8 60.6 58.1 59.7 58.2 37.7 37.5 22.8 26.1 56.7 51.1
Trans-Google 74.7 72.7 90.3 83.2 62.4 59.1 68.2 62.0 42.5 48.3 30.6 28.9 61.4 59.0
Trans-NLLB 65.6 54.1 85.7 78.2 60.5 58.2 68.4 63.4 35.4 43.6 28.4 27.7 57.3 54.2
bloomz-7b1 Native-Basic 1.6 0.9 36.5 18.9 3.7 11.8 - - 7.1 10.5 - - 12.2 10.5
EN-Basic 1.9 2.2 67.5 55.2 48.2 40.7 - - 11.8 6.5 - - 32.4 26.2
Native-CoT 1.0 1.4 37.9 17.3 1.2 13.5 - - 5.2 11.1 - - 11.3 10.8
EN-CoT 1.7 1.6 61.3 52.8 37.6 34.7 - - 10.0 6.9 - - 27.7 24.0
XLT 1.9 1.5 58.6 49.2 35.4 35.3 - - 8.6 5.9 - - 26.1 23.0
Trans-Google 2.5 3.0 67.5 62.8 44.4 44.2 - - 15.6 23.0 - - 32.5 33.2
Trans-NLLB 2.0 2.9 64.3 61.2 44.1 43.6 - - 12.8 21.3 - - 30.8 32.2
Mistral-7B-Instruct Native-Basic 15.5 4.9 69.7 50.0 50.6 37.0 44.6 44.8 7.8 8.1 26.3 24.4 35.7 28.2
EN-Basic 33.7 8.8 42.5 33.8 55.5 46.2 47.0 46.6 6.8 8.0 21.7 21.1 34.5 27.4
Native-CoT 23.1 8.0 67.7 49.9 50.2 38.3 44.3 44.2 7.7 8.2 25.5 21.1 36.4 28.3
EN-CoT 37.3 13.1 50.9 38.9 54.2 46.8 46.6 46.4 11.3 12.0 18.7 18.8 36.5 29.3
XLT 43.0 15.0 78.3 57.9 48.4 44.3 47.9 47.2 9.4 10.4 17.1 19.6 40.7 32.4
Trans-Google 42.6 39.4 67.0 57.5 56.4 53.9 51.4 52.0 16.3 19.7 31.9 36.5 44.3 43.2
Trans-NLLB 32.3 30.8 62.1 52.3 54.4 51.9 52.2 53.6 14.5 19.3 31.0 37.3 41.1 40.9
Llama-2-13b-Chat Native-Basic 22.7 4.9 59.5 48.4 39.9 33.7 55.2 48.2 20.7 9.6 28.4 23.8 37.7 28.1
EN-Basic 28.7 4.4 63.9 51.6 48.2 39.8 59.6 56.8 20.9 17.8 31.3 30.2 42.1 33.4
Native-CoT 26.9 4.9 59.0 49.3 38.6 33.5 56.2 47.8 17.9 7.8 28.4 22.7 37.8 27.7
EN-CoT 29.5 5.5 68.2 51.0 46.2 41.8 57.8 56.6 20.5 17.3 30.7 28.0 42.1 33.4
XLT 32.8 6.5 68.1 52.7 56.9 47.3 56.0 54.2 19.6 16.8 22.0 18.1 42.6 32.6
Trans-Google 38.4 40.1 77.8 70.4 46.1 46.1 59.2 54.6 32.6 37.8 35.1 38.0 48.2 47.8
Trans-NLLB 32.8 30.4 72.7 67.1 45.6 45.2 58.1 56.2 26.7 34.7 33.4 37.3 44.9 45.1
Llama-2-70B-Chat Native-Basic 35.7 5.6 64.2 48.0 43.0 36.0 53.3 50.4 28.9 10.4 30.1 26.8 42.5 29.5
EN-Basic 42.5 7.7 70.7 52.0 52.7 41.9 61.9 52.8 25.7 21.5 30.2 35.3 47.3 35.2
Native-CoT 35.5 5.6 65.3 46.8 41.0 35.6 56.0 49.6 25.3 9.9 26.0 25.2 41.5 28.8
EN-CoT 45.6 7.0 80.7 56.3 52.7 40.9 66.5 57.0 32.7 25.7 29.8 32.0 51.3 36.5
XLT 49.0 8.4 76.4 54.7 57.3 48.4 56.6 51.6 26.5 26.7 19.3 11.5 47.5 33.6
Trans-Google 55.5 50.0 86.3 79.7 55.3 53.0 69.4 64.2 38.7 43.1 33.1 36.7 56.4 54.4
Trans-NLLB 46.5 39.7 83.3 75.6 53.7 51.0 70.5 62.4 17.8 24.7 32.4 36.2 50.7 48.3
Table 4: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting.
Model Prompt type de ru fr zh es ja sw th bn te avg
ChatGPT Native-Basic 48.8 42.8 42.8 36.0 50.0 46.0 30.8 21.6 15.6 9.6 34.4
EN-Basic 49.2 56.0 48.4 52.4 57.2 38.4 42.0 27.2 28.8 11.2 41.1
Native-CoT 66.0 69.6 62.4 64.4 70.0 58.0 49.2 28.4 20.8 10.0 49.9
EN-CoT 74.8 72.4 71.2 67.2 75.2 62.0 58.0 51.6 52.8 26.0 61.1
XLT 70.8 73.6 69.6 68.8 72.8 66.8 65.6 56.8 50.8 27.2 62.3
Trans-Google 76.8 76.4 75.2 73.2 76.0 70.4 73.6 76.0 74.0 67.2 73.9
Trans-NLLB 70.0 63.2 71.2 58.4 71.6 59.2 61.2 44.4 55.6 55.2 61.0
bloomz-7b1 Native-Basic 1.2 1.2 2.0 2.8 1.6 0.8 0.8 0.0 1.6 1.2 1.3
EN-Basic 2.0 1.6 2.4 2.8 1.6 1.2 2.0 1.2 3.6 2.0 2.0
Native-CoT 0.0 0.4 1.2 1.6 1.6 1.2 2.4 0.4 1.2 1.6 1.2
EN-CoT 2.0 1.2 2.4 2.0 0.8 2.0 1.6 1.2 2.0 1.6 1.7
XLT 0.8 1.2 2.0 3.2 1.6 2.4 2.0 0.8 0.8 2.4 1.7
Trans-Google 3.2 2.0 2.4 2.4 2.4 2.4 2.0 3.2 2.0 4.8 2.7
Trans-NLLB 2.4 1.6 3.2 0.8 2.0 2.0 3.6 2.4 2.4 3.2 2.4
Mistral-7B-Instruct Native-Basic 7.6 14.4 12.0 19.2 30.8 8.8 4.0 4.4 6.8 4.4 11.2
EN-Basic 38.4 36.4 31.6 28.0 42.4 25.6 7.6 9.6 16.0 2.0 23.8
Native-CoT 9.6 24.0 16.8 26.8 38.4 22.8 6.0 7.6 17.2 1.2 17.0
EN-CoT 39.2 42.0 36.0 33.6 42.0 30.8 8.0 21.6 18.4 4.4 27.6
XLT 43.6 51.6 45.2 38.4 45.2 34.0 10.4 23.6 19.6 6.4 31.8
Trans-Google 42.0 46.8 41.2 44.0 42.0 39.6 38.8 35.6 42.0 41.2 41.3
Trans-NLLB 37.6 30.0 34.0 24.8 38.0 29.6 31.6 26.4 31.2 34.0 31.7
Llama-2-13b-Chat Native-Basic 25.2 20.0 25.6 24.4 22.0 18.8 3.6 7.2 5.2 3.6 15.6
EN-Basic 32.4 26.4 32.0 26.0 34.8 20.8 3.2 5.6 5.6 3.2 19.0
Native-CoT 29.2 23.6 29.2 27.6 28.4 23.2 2.8 7.2 6.4 3.2 18.1
EN-CoT 34.0 32.4 32.0 24.4 35.6 18.4 5.6 6.8 6.0 3.6 19.9
XLT 34.4 34.4 33.6 29.6 37.2 27.6 4.8 8.4 9.2 3.6 22.3
Trans-Google 38.0 40.4 36.8 35.6 44.8 34.8 38.4 39.2 42.8 40.0 39.1
Trans-NLLB 29.6 33.2 38.8 31.2 28.0 36.0 32.0 24.8 35.6 29.2 31.8
Llama-2-70B-Chat Native-Basic 34.8 28.4 38.8 38.8 41.2 32.0 4.4 8.4 7.6 2.0 23.6
EN-Basic 50.4 39.2 48.0 40.0 48.0 29.6 6.0 8.8 11.6 4.4 28.6
Native-CoT 41.2 31.6 36.4 35.6 36.8 31.2 6.4 5.2 9.2 1.6 23.5
EN-CoT 49.6 48.0 50.0 38.0 48.4 39.6 7.6 7.2 10.4 2.8 30.2
XLT 52.0 49.6 49.6 47.2 52.0 43.6 8.0 8.0 15.6 2.0 32.8
Trans-Google 56.8 56.4 54.4 54.8 56.4 54.0 51.6 46.0 51.6 50.8 53.3
Trans-NLLB 49.6 43.6 49.2 41.2 50.4 45.2 43.6 32.0 42.0 41.2 43.8
Table 5: Accuracy scores across various languages on the MGSM benchmark.
Model Prompt type zh it vi tr id sw th et ta ht qu avg
ChatGPT Native-Basic 88.0 91.8 74.0 81.4 85.4 77.2 65.2 85.4 49.6 63.4 50.0 72.3
EN-Basic 90.0 89.8 85.0 86.0 87.2 78.2 75.0 81.4 58.2 65.8 54.8 76.1
Native-CoT 87.0 92.6 72.8 80.8 83.8 75.4 66.8 84.8 48.6 63.2 55.2 72.4
EN-CoT 90.4 92.2 87.0 89.6 90.2 85.6 74.8 85.8 61.4 69.2 50.2 78.6
XLT 89.4 91.2 87.4 88.0 88.8 82.4 76.4 91.0 60.6 76.8 50.4 79.3
Trans-Google 90.8 91.6 88.4 85.8 88.8 79.4 82.6 88.2 85.6 81.6 73.2 84.5
Trans-NLLB 85.6 89.2 82.4 85.8 87.0 81.4 73.8 85.4 80.6 76.2 55.6 79.7
bloomz-7b1 Native-Basic 46.6 48.6 14.4 1.6 48.4 39.0 20.0 0.0 19.0 2.8 20.6 21.4
EN-Basic 78.2 55.6 68.6 50.2 62.8 56.8 49.6 50.0 71.4 50.0 50.4 56.5
Native-CoT 43.4 50.0 20.2 0.6 48.6 23.0 39.2 0.0 17.6 0.0 9.4 20.9
EN-CoT 67.4 53.4 63.0 50.4 57.4 51.4 49.6 49.4 64.0 49.6 50.6 53.9
XLT 63.8 49.6 62.4 45.6 64.0 49.0 51.2 46.0 52.8 48.0 36.6 50.5
Trans-Google 68.0 68.6 66.0 65.2 68.8 60.4 59.4 67.2 61.8 61.6 57.6 63.7
Trans-NLLB 64.0 67.2 61.6 63.6 64.6 62.2 57.4 62.8 62.8 61.6 54.2 61.8
Mistral-7B-Instruct Native-Basic 67.2 82.2 59.8 55.0 65.0 47.6 51.8 36.6 49.2 51.2 43.6 54.2
EN-Basic 48.6 43.6 35.4 30.6 43.6 37.8 39.8 28.6 35.2 29.4 25.0 34.9
Native-CoT 64.0 80.4 58.6 54.6 65.4 45.4 50.0 40.0 44.2 51.2 48.2 53.8
EN-CoT 55.8 52.2 44.6 43.8 52.2 39.8 46.0 32.6 29.2 39.4 28.2 40.8
XLT 82.6 81.4 70.8 66.8 77.8 47.8 64.2 53.6 52.0 56.6 44.0 61.5
Trans-Google 69.4 64.8 66.8 61.0 68.8 52.2 62.0 60.8 59.8 52.0 43.6 59.2
Trans-NLLB 60.8 63.4 62.2 59.2 63.0 50.8 51.4 60.6 55.0 51.0 27.4 54.4
Llama-2-13b-Chat Native-Basic 65.0 62.2 51.4 50.4 57.6 46.2 48.4 50.0 40.2 47.2 47.0 50.1
EN-Basic 61.2 74.2 56.2 52.8 62.0 52.0 50.6 50.6 50.2 46.4 48.4 54.3
Native-CoT 62.8 64.6 49.6 53.8 64.8 49.8 51.8 45.4 32.6 49.8 46.6 50.9
EN-CoT 67.4 71.8 65.4 51.4 68.2 48.2 49.0 46.8 48.6 50.4 45.6 54.5
XLT 65.4 72.6 66.2 57.2 70.0 47.0 49.2 50.8 50.2 50.6 46.6 56.0
Trans-Google 77.8 80.4 75.2 75.0 76.4 66.6 67.6 74.0 71.8 68.8 63.2 71.9
Trans-NLLB 73.0 75.6 69.6 74.4 73.2 67.4 62.4 73.8 66.2 68.0 51.2 68.2
Llama-2-70B-Chat Native-Basic 61.6 81.6 49.4 49.4 55.4 50.6 46.8 49.8 41.0 46.4 44.6 51.5
EN-Basic 74.6 79.4 58.0 53.6 63.2 48.8 50.2 49.4 50.4 49.0 51.0 55.3
Native-CoT 65.8 78.0 52.2 51.8 54.8 49.2 49.2 50.2 40.0 43.2 36.2 50.5
EN-CoT 80.4 88.0 73.6 65.4 77.8 53.0 50.0 56.0 48.0 49.8 50.6 61.2
XLT 79.8 82.0 67.4 64.6 74.4 49.8 51.8 55.0 47.8 46.2 48.2 58.7
Trans-Google 87.2 88.0 83.6 82.2 89.4 76.6 77.6 83.4 83.6 76.4 68.4 80.9
Trans-NLLB 83.2 86.6 80.2 79.8 85.8 74.4 71.4 79.2 79.6 76.2 58.2 77.1
Table 6: Accuracy scores across various languages on the XCOPA benchmark.
Model Prompt type de ru fr zh es vi tr sw ar el th bg hi ur avg
ChatGPT Native-Basic 59.0 58.8 60.2 54.0 60.2 49.2 51.6 51.0 50.6 58.0 39.6 54.8 42.8 40.4 52.2
EN-Basic 68.6 58.2 67.4 62.2 68.4 63.0 65.6 65.2 62.4 64.6 56.4 65.4 55.8 59.0 63.0
Native-CoT 59.4 54.2 58.0 51.8 58.6 47.6 53.0 50.8 51.2 54.6 37.2 54.4 40.2 37.4 50.6
EN-CoT 62.6 56.4 61.4 57.4 65.8 57.6 58.0 54.0 53.4 59.0 51.0 59.8 48.6 45.0 56.4
XLT 63.0 57.8 61.4 58.4 63.4 59.8 61.4 58.0 57.8 60.4 55.0 59.6 53.2 59.2 59.2
Trans-Google 65.6 59.6 65.2 62.6 62.6 58.6 60.4 57.6 63.2 62.2 56.4 60.0 57.0 55.8 60.5
Trans-NLLB 63.4 62.2 61.6 57.4 62.8 55.6 59.4 58.8 62.4 63.4 54.2 61.6 52.8 53.0 59.2
bloomz-7b1 Native-Basic 0.4 13.4 0.2 6.6 1.4 0.0 6.8 18.2 1.6 5.2 26.6 15.4 17.8 2.8 8.3
EN-Basic 39.8 42.8 50.8 52.4 52.2 51.4 34.2 42.4 45.6 37.2 33.8 40.4 49.2 43.0 43.9
Native-CoT 0.4 3.0 1.2 1.2 1.2 0.2 9.0 27.2 1.6 0.8 33.4 12.4 20.0 3.8 8.2
EN-CoT 36.2 35.2 37.4 42.2 37.4 37.2 33.2 34.8 36.2 33.6 33.2 34.2 37.6 34.4 35.9
XLT 38.2 34.4 35.0 34.0 35.0 36.0 37.4 35.4 34.6 35.6 35.0 36.6 33.8 34.0 35.4
Trans-Google 45.0 43.4 44.2 44.0 45.2 44.8 43.8 44.0 44.0 44.6 44.4 44.8 43.4 44.4 44.3
Trans-NLLB 45.6 43.0 44.0 44.0 45.4 42.4 43.6 43.4 44.6 44.6 43.2 44.8 42.8 42.0 43.8
Mistral-7B-Instruct Native-Basic 50.4 55.6 59.2 46.0 59.0 33.4 38.8 33.0 34.2 34.2 39.2 46.6 37.0 33.2 42.8
EN-Basic 56.4 54.6 59.8 54.0 56.8 51.4 46.8 37.6 45.8 49.4 47.0 54.4 46.4 41.8 50.2
Native-CoT 50.0 55.0 58.4 47.6 54.6 35.8 38.2 32.2 37.6 35.4 40.0 52.0 36.8 33.8 43.4
EN-CoT 55.0 52.2 58.0 52.4 57.0 50.4 48.0 38.0 48.6 51.2 45.8 54.2 46.8 42.0 50.0
XLT 48.2 44.6 49.6 49.4 52.4 46.0 48.0 39.0 42.2 46.4 45.4 46.6 44.0 42.6 46.0
Trans-Google 58.6 54.2 59.2 52.6 59.0 55.0 54.6 53.0 56.4 58.2 48.8 56.8 52.4 50.6 55.0
Trans-NLLB 57.0 52.4 55.8 50.2 58.2 53.0 54.2 49.4 53.0 56.4 47.4 55.2 50.0 49.6 53.0
Llama-2-13b-Chat Native-Basic 41.4 40.2 44.0 38.6 42.8 32.4 34.6 31.6 32.8 34.2 34.0 37.4 31.4 33.6 36.4
EN-Basic 50.2 47.4 51.6 45.0 51.8 43.0 41.8 37.8 38.8 42.6 36.4 45.0 38.4 37.8 43.4
Native-CoT 39.4 42.0 43.4 32.6 42.6 31.8 31.4 33.4 31.2 35.2 32.8 38.2 32.6 33.2 35.7
EN-CoT 45.6 46.8 48.8 44.4 46.6 44.8 41.8 38.6 43.2 43.4 38.6 46.2 42.0 40.8 43.7
XLT 59.6 55.8 56.4 54.0 59.8 55.6 48.2 37.8 49.4 49.0 44.4 52.0 48.4 49.2 51.4
Trans-Google 50.4 44.2 45.4 44.6 46.0 46.0 47.6 42.8 48.4 48.2 43.4 45.4 45.4 47.4 46.1
Trans-NLLB 48.6 46.6 47.0 43.2 44.6 43.6 49.0 43.2 44.0 46.0 41.6 48.6 45.6 43.4 45.4
Llama-2-70B-Chat Native-Basic 44.0 42.0 45.4 42.6 45.6 38.4 38.4 32.6 35.0 37.6 33.0 41.8 34.8 34.8 39.0
EN-Basic 53.6 54.6 57.0 49.6 55.6 46.0 42.8 32.4 50.2 46.2 38.6 52.4 37.6 34.8 46.5
Native-CoT 40.4 42.2 45.4 38.4 41.4 38.4 36.6 32.8 35.2 37.4 32.6 41.0 33.2 36.2 37.9
EN-CoT 53.6 52.8 56.4 50.4 56.8 46.0 40.6 33.4 44.6 47.8 38.2 48.2 37.6 36.6 45.9
XLT 56.0 59.4 59.6 55.2 61.2 52.6 51.4 36.4 44.4 55.4 44.6 57.8 51.2 45.8 52.2
Trans-Google 58.8 53.4 56.8 56.4 54.8 51.8 55.4 49.6 57.2 56.4 50.2 57.4 50.8 46.6 54.0
Trans-NLLB 56.4 52.8 54.6 49.8 58.6 50.2 53.4 51.0 52.0 56.0 48.8 52.4 49.0 45.6 52.2
Table 7: Accuracy scores across various languages on the XNLI benchmark.
Model Prompt type de fr zh es ja ko avg
ChatGPT Native-Basic 62.0 53.6 46.6 46.6 49.0 40.6 49.7
EN-Basic 67.6 68.0 58.8 71.4 55.8 50.4 62.0
Native-CoT 61.8 55.0 48.6 48.8 44.0 43.4 50.3
EN-CoT 67.6 64.0 61.2 70.0 55.8 51.2 61.6
XLT 57.4 63.8 59.8 59.2 58.2 58.2 59.4
Trans-Google 69.0 69.6 66.0 71.4 65.0 62.0 67.2
Trans-NLLB 67.0 70.6 68.6 70.2 65.4 63.4 67.5
Mistral-7B-Instruct Native-Basic 40.6 47.0 49.2 44.2 41.8 44.8 44.6
EN-Basic 46.8 47.8 47.8 46.8 45.8 46.6 46.9
Native-CoT 43.8 50.2 38.8 43.6 45.0 44.2 44.3
EN-CoT 46.2 47.4 47.8 47.0 44.8 46.4 46.6
XLT 47.4 49.6 47.6 46.6 48.2 47.2 47.8
Trans-Google 51.2 49.8 54.0 49.6 52.4 52.0 51.5
Trans-NLLB 50.6 52.8 52.4 50.8 54.2 53.6 52.4
Llama-2-13b-Chat Native-Basic 50.8 57.2 54.0 58.0 55.8 48.2 54.0
EN-Basic 60.2 61.0 58.6 59.8 58.2 56.8 59.1
Native-CoT 50.4 58.8 59.0 55.8 56.8 47.8 54.8
EN-CoT 59.2 55.8 58.6 59.2 56.4 56.6 57.6
XLT 54.8 58.0 53.6 56.6 56.8 54.2 55.7
Trans-Google 56.6 62.0 59.6 61.6 56.2 54.6 58.4
Trans-NLLB 56.2 60.0 57.4 59.4 57.6 56.2 57.8
Llama-2-70B-Chat Native-Basic 53.4 49.8 55.6 61.0 46.8 50.4 52.8
EN-Basic 62.8 66.2 58.4 67.0 55.2 52.8 60.4
Native-CoT 53.0 53.4 53.6 65.4 54.6 49.6 54.9
EN-CoT 65.0 70.8 65.0 70.2 61.6 57.0 64.9
XLT 57.0 61.6 57.6 57.2 49.4 51.6 55.7
Trans-Google 70.6 70.6 68.0 72.2 65.6 64.2 68.5
Trans-NLLB 69.8 73.4 69.4 71.2 68.8 62.4 69.2
Table 8: Accuracy scores across various languages on the PAWS-X benchmark.
Model Prompt type de ru fr zh es ja vi tr th avg
ChatGPT Native-Basic 44.1 30.5 46.4 31.4 40.2 20.2 33.0 39.2 33.6 35.4
EN-Basic 36.9 30.5 43.3 28.9 44.1 43.5 34.7 32.7 34.0 36.5
Native-CoT 43.6 22.2 46.1 30.0 38.3 33.9 34.1 38.3 31.9 35.4
EN-CoT 44.6 37.4 49.7 38.5 48.0 52.4 32.6 42.0 40.5 42.9
XLT 36.6 31.0 39.3 31.8 44.0 43.6 37.3 37.9 37.2 37.6
ransg 42.0 39.2 42.7 48.6 40.8 46.4 37.8 44.2 52.3 43.8
Trans-NLLB 39.2 34.6 26.7 31.6 29.1 45.3 41.2 41.2 45.9 37.2
bloomz-7b1 Native-Basic 0.6 3.0 7.6 12.1 11.2 7.5 7.6 0.0 20.9 7.8
EN-Basic 7.5 3.7 12.3 21.4 12.2 12.3 13.3 2.1 11.0 10.6
Native-CoT 0.2 0.9 5.9 8.6 8.3 6.0 6.7 0.0 22.2 6.5
EN-CoT 4.0 3.0 11.4 17.9 13.9 8.7 11.1 1.7 12.2 9.3
XLT 5.7 2.8 10.2 14.8 10.1 7.1 9.6 1.4 10.4 8.0
Trans-Google 13.5 11.5 10.7 25.7 12.5 22.5 12.8 11.7 34.2 17.2
Trans-NLLB 11.7 8.7 7.2 15.2 9.3 24.5 13.1 11.2 31.3 14.7
Mistral-7B-Instruct Native-Basic 8.5 5.2 8.7 7.2 9.5 7.4 8.0 2.6 13.5 7.8
EN-Basic 7.9 5.0 7.5 5.1 8.7 6.7 6.3 5.3 10.6 7.0
Native-CoT 9.1 5.4 7.7 8.1 8.2 7.9 7.3 2.8 13.6 7.8
EN-CoT 11.2 7.8 16.0 8.4 14.9 13.1 7.9 7.6 16.4 11.5
XLT 9.7 7.2 10.4 8.4 10.4 10.5 9.2 6.6 14.2 9.6
Trans-Google 14.6 13.8 14.9 17.7 17.0 22.5 13.4 15.1 24.4 17.0
Trans-NLLB 13.3 12.7 10.5 14.9 11.8 24.1 13.8 13.5 25.2 15.5
Llama-2-13b-Chat Native-Basic 15.0 13.6 31.3 20.6 29.7 13.8 21.2 5.8 13.4 18.3
EN-Basic 28.5 11.6 28.7 13.9 27.2 21.0 15.3 15.6 20.0 20.2
Native-CoT 14.6 10.4 29.1 13.3 23.6 10.5 23.8 5.6 10.1 15.7
EN-CoT 28.2 12.6 31.1 11.9 28.9 15.3 15.4 18.3 16.3 19.8
XLT 23.6 17.0 27.5 10.3 26.2 18.2 14.7 16.4 17.2 19.0
Trans-Google 31.1 29.9 34.6 35.1 31.7 35.4 30.8 31.7 43.9 33.8
Trans-NLLB 26.1 26.6 19.8 27.4 18.5 36.2 32.1 29.2 40.2 28.4
Llama-2-70B-Chat Native-Basic 36.7 23.8 35.2 15.9 39.3 24.7 26.7 8.6 12.1 24.8
EN-Basic 33.2 18.1 32.9 18.8 33.7 26.6 16.3 20.7 22.3 24.7
Native-CoT 34.8 19.5 33.9 13.1 38.5 13.1 24.1 9.2 10.6 21.9
EN-CoT 39.5 24.6 39.0 24.2 41.0 35.2 25.3 26.4 25.0 31.1
XLT 29.8 22.4 29.6 18.0 31.3 29.5 25.0 27.3 26.1 26.6
Trans-Google 37.3 34.0 37.1 43.5 35.4 48.0 35.8 38.3 47.9 39.7
Trans-NLLB 16.7 16.4 11.9 18.5 14.9 26.8 19.7 21.8 27.6 19.4
Table 9: F1 scores across various languages on the MKQA benchmark.
Model Prompt type fr zh es vi tr avg
ChatGPT Native-Basic 29.2 39.3 26.9 34.4 29.9 31.9
EN-Basic 28.9 38.8 27.8 37.9 30.0 32.7
Native-CoT 28.8 38.5 26.1 34.0 27.9 31.1
EN-CoT 25.4 35.1 26.0 33.5 28.6 29.7
XLT 24.2 25.5 18.1 23.4 26.1 23.4
Trans-Google 27.2 36.2 26.3 32.6 28.9 30.3
Trans-NLLB 26.4 29.7 26.1 31.5 27.7 28.3
bloomz-7b1 Native-Basic 14.6 24.3 20.0 7.7 8.2 14.9
EN-Basic 20.1 23.9 20.9 20.6 14.2 19.9
Native-CoT 18.2 25.4 24.1 1.7 8.0 15.5
EN-CoT 18.0 26.1 21.6 19.3 11.3 19.3
XLT 12.2 19.9 19.3 14.5 5.3 14.2
Trans-Google 10.0 14.2 12.1 9.0 10.7 11.2
Trans-NLLB 10.5 8.6 12.5 9.7 11.5 10.6
Mistral-7B-Instruct Native-Basic 23.0 34.0 22.3 25.8 24.4 25.9
EN-Basic 20.9 16.5 21.5 28.0 21.1 21.6
Native-CoT 19.7 33.6 22.1 26.4 21.1 24.6
EN-CoT 20.6 12.1 19.9 22.2 18.8 18.7
XLT 15.4 16.5 14.7 21.7 19.6 17.6
Trans-Google 26.8 34.9 26.4 39.5 36.5 32.8
Trans-NLLB 26.8 30.0 26.6 40.6 37.3 32.2
Llama-2-13b-Chat Native-Basic 27.7 21.9 25.3 38.8 23.8 27.5
EN-Basic 25.7 38.2 23.6 37.7 30.2 31.1
Native-CoT 27.9 29.0 24.8 31.8 22.7 27.2
EN-CoT 24.0 39.4 23.1 36.4 28.0 30.2
XLT 24.2 17.7 22.4 23.6 18.1 21.2
Trans-Google 28.0 42.9 27.9 41.6 38.0 35.7
Trans-NLLB 27.5 37.5 26.9 41.6 37.3 34.2
Llama-2-70B-Chat Native-Basic 28.8 34.5 27.3 29.7 26.8 29.4
EN-Basic 29.0 31.8 24.3 35.7 35.3 31.2
Native-CoT 25.3 29.5 26.7 22.4 25.2 25.8
EN-CoT 27.0 35.2 22.1 34.8 32.0 30.2
XLT 18.1 29.7 15.2 14.2 11.5 17.7
Trans-Google 26.8 39.7 27.1 38.7 36.7 33.8
Trans-NLLB 26.6 37.5 26.3 39.0 36.2 33.1
Table 10: ROUGE-1 scores across various languages on the XL-sum benchmark.
Figure 6: The LLM-as-a-judge prompt for GPT-4-Turbo.
Figure 7: Prompt to check whether answering a request needs local cultural knowledge.

### v2

@@
 Abstract
-Large language models (LLMs) have demonstrated strong multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora.
-Existing works leverage this phenomenon to improve their multilingual performances on NLP tasks.
-In this work, we extend the evaluation from NLP tasks to real user queries.
-We find that even though translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios.
-For culture-related tasks that need deep language understanding, prompting in the native language proves to be more promising since it can capture the nuances related to culture and language.
-Therefore, we advocate for more efforts towards the development of strong multilingual LLMs instead of just English-centric LLMs.
+Large language models (LLMs) have demonstrated multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora.
+Existing works leverage this phenomenon to improve their multilingual performances through translation, primarily on natural language processing (NLP) tasks.
+This work extends the evaluation from NLP tasks to real user queries and from English-centric LLMs to non-English-centric LLMs.
+While translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios.
+For culture-related tasks that need deep language understanding, prompting in the native language tends to be more promising as it better captures the nuances of culture and language.
+Our experiments reveal varied behaviors among different LLMs and tasks in the multilingual context.
+Therefore, we advocate for more comprehensive multilingual evaluation and more efforts toward developing multilingual LLMs beyond English-centric ones.
 Is Translation All You Need?
 A Study on Solving Multilingual Tasks with Large Language Models
-[…]
+Chaoqun Liuthanks: Chaoqun Liu is under the Joint PhD Program between DAMO Academy and Nanyang Technological University.
+12  Wenxuan Zhangthanks: Wenxuan Zhang is the corresponding author.
+23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bing23 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore {chaoqun.liu,saike.zwx,zhaoyiran.zyr,l.bing}@alibaba-inc.com anhtuan.luu@ntu.edu.sg
 1 Introduction
@@
 (2023b).
-Figure 1: Illustration of two kinds of LLMs on different tasks.
+Figure 1: Illustration of two types of LLMs on tasks with varying language dependencies.
 "English-centric LLMs" refers to LLMs trained mainly in English corpora.
-"Multilingual LLMs" refers to ideal LLMs that are equally capable in all languages.
+"Multilingual LLMs" refers to ideal LLMs equally capable in all languages.
 To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al.
 (2018); Ponti et al.
+(2020); Artetxe et al.
+(2023); Moghe et al.
 (2023).
-Recent investigations have expanded this concept by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al.
+Recent investigations have expanded this idea by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al.
 (2023); Qin et al.
-(2023); Etxaniz et al.
+(2023b); Etxaniz et al.
 (2023) for multilingual NLP tasks.
 For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a).
-Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then proceed to solve problems step by step, leads to better performance.
-Figure 2: Examples illustrating how translation can both improve (top) and degrade (bottom) the performance of LLMs.
-The example in Chinese is drawn from MGSM and the example in Swahili is drawn from M3Exam, respectively.
-Translation is beneficial when the questions are language-independent.
-However, for questions requiring deep cultural knowledge, translation can hinder the ability to answer correctly.
-Despite these advancements, the methodologies employed across different studies vary considerably, and the specific impact of translation in improving multilingual task performance has not been rigorously analyzed.
-Furthermore, the focus of these studies tends to be on specific NLP tasks and English-centric LLMs, with less attention given to real-world scenarios of multilingual user requests.
-This gap highlights a need for more nuanced research into the effectiveness of translation techniques across a spectrum of multilingual contexts.
-In this study, we conduct a thorough investigation into the utility of translation with large language models for a variety of scenarios, including multilingual NLP tasks and real-world multilingual user queries.
-We compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages against several baselines, such as using native language prompts and cross-lingual prompts Huang et al.
+Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then solve them step by step, improves performance.
+Figure 2: Examples illustrating how translation can both improve (a) and degrade (b) the performance of LLMs.
+The Chinese example is from MGSM Shi et al.
+(2022) and the Swahili example is from M3Exam Zhang et al.
+(2023a).
+Translation is beneficial when the questions are semantically equivalent across languages.
+However, for questions that demand deep cultural knowledge, translation can hinder the ability to answer accurately.
+Despite these advancements, methodologies in various studies differ significantly, and the impact of translation on multilingual task performance remains underexplored.
+Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, with less attention given to real-world user requests in various languages.
+This gap highlights a need for more nuanced research into the effectiveness of translation techniques across multilingual contexts.
+As shown in Figure 1, we hypothesize that English-centric LLMs generally perform better with English translations of prompts, while "Multilingual LLMs" excel with native prompts, particularly for tasks highly dependent on language.
+In this study, we perform an in-depth analysis of the utility of translation with large language models for various scenarios.
+Firstly, we compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages (i.e., the “translate-test” method), against several baselines on multilingual NLP tasks.
+Secondly, we extend the evaluation to real user queries, which are more likely to contain knowledge related to culture and language.
+Thirdly, We broaden the scope of LLM evaluations to include non-English-centric models to explore how they differ in behavior from English-centric LLMs.
+Our results demonstrate that simply translating queries into English can already achieve the best results in multiple NLP task categories.
+For real user queries, the effect of translation depends on the languages and the LLMs.
+When working with advanced LLMs and certain languages, employing prompts in native languages appears to be the more effective strategy.
+In addition, the non-English-centric LLMs also behave differently from English-centric LLMs, where prompts in the native languages yield superior results by capturing the nuances related to culture and language.
+The main contributions of this work can be summarized as follows:
+  • We compare various multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs.
+Our study also identifies factors that affect performance in multilingual tasks.
+  • We broaden our multilingual evaluation to encompass actual user queries and LLMs that are not primarily centered on English, thereby addressing the limitations of previous works.
+  • Our study reveals the shortcomings of current multilingual evaluations, underscoring the need for more comprehensive assessments that broaden the range of benchmarks and diversify the types of LLMs considered.
+2 Translation for NLP Tasks
+This section explores various prompting strategies across multiple languages and LLMs, covering a wide range of NLP tasks.
+This helps us understand how different prompting methods and other factors affect task performance.
+2.1 Experiment Setup
+2.1.1 Tasks
+We conduct assessments on six benchmarks covering reasoning, understanding, and generation tasks that encapsulate various abilities of LLMs: MGSM Shi et al.
+(2022), XCOPA Ponti et al.
+(2020), XNLI Conneau et al.
+(2018), PAWS-X (Yang et al., 2019), MKQA (Longpre et al., 2021) and XL-Sum Hasan et al.
 (2021).
-Our experiments also span across a diverse set of LLMs and translation tools to ensure the broad applicability of our findings.
-Our results demonstrate that simply translating queries into English can already achieve the best results in multiple task categories.
-Subsequently, we investigate the utility of translation in real-world scenarios by conducting tests with requests from ShareGPT [^1] https://sharegpt.com/.
-We discover that translation tends to yield better responses for low-resource languages.
-However, when working with high-resource languages and advanced LLMs, employing prompts in native languages appears to be the more effective strategy.
-To further analyze such behaviors, we extend our evaluations to LLMs optimized for languages other than English, focusing specifically on culture-related tasks.
-We observe that for these models, prompts in the native languages yield superior results by capturing the nuances related to culture and language, as illustrated in Figure 2 (b).
-In summary, as illustrated in Figure 1, while translation can improve the performance of English-centric LLMs, it may not be optimal for all scenarios.
-In this work, we highlight the need for robust LLMs that are equally proficient across all languages to effectively address a broader range of cases.
-Therefore, we advocate for more efforts towards the development of multilingual LLMs.
-2 Translation for NLP Tasks
-In this section, we explore various prompting strategies across multiple languages and LLMs, encompassing a wide spectrum of NLP tasks.
-This allows us to understand how different prompting methods impact the performance of tasks that theoretically do not depend on any specific language.
-2.1 Experiment Setup
-Original Question
-制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
-Native-Basic
-{Original Question }
-您的最终答案的格式应为:"答案: <阿拉伯数字>".
-EN-Basic
-{Original Question }
-You should format your final answer as "Answer: <Arabic numeral>".
-Native-CoT
-{Original Question }
-让我们一步步思考。
-您的最终答案的格式应为:"答案: <阿拉伯数字>".
-EN-CoT
-{Original Question }
-Let’s think step by step in English.
-You should format your final answer as "Answer: <Arabic numeral>".
-XLT
-I want you to act as an arithmetic reasoning expert for Chinese.
-Request: {Original Question}
-You should retell the request in English.
-You should do step-by-step answer to obtain a number answer.
-You should step-by-step answer the request.
-You should tell me the answer in this format ’Answer :’.
-Trans-Google
-Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
-Let’s think step by step.
-You should format your final answer as "Answer: <Arabic numeral>".
-Trans-NLLB
-To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
-Let’s think step by step.
-You should format your final answer as "Answer: <Arabic numeral>".
-Table 1: An example of zero-shot prompts for a Chinese problem.
-For Native-Basic, EN-Basic, Native-CoT, EN-CoT and XLT, we provide the original Chinese question as input and expect an answer in the corresponding format; for Trans-Google and Trans-NLLB, we input the translated question in English, and expect a step-by-step solution in English.
-To obtain the desirable output format, we instruct the models to output in specific format.
-We perform assessments on six benchmarks on reasoning, understanding, and generation tasks that can encapsulate various abilities of LLMs.
-We follow the task and language selection from Huang et al.
-(2023) to consider a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu).
+Following Huang et al., 2023, we choose a subset of 9 languages for MKQA and 5 languages for XL-Sum.
+For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks.
+More details of the benchmarks can be found in Appendix A.1.
+These tasks cover a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu).
 We categorize languages larger than 1% frequency in Common Crawl [^1] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages as high-resource languages (i.e., de, ru, fr, zh, es, ja, it and vi), and the rest as low-resource languages.
 We exclude English since we want to evaluate the efficient prompting strategy for non-English tasks.
-As our prompting strategies involve methods with translators, we do not test on the FLORES-200 dataset.
-2.1.1 Tasks
-We evaluate the models and prompting strategies on the following tasks: 1) the MGSM benchmark Shi et al.
-(2022), which tests models on grade school math problems.
-2) the XCOPA benchmark Ponti et al.
-(2020), which presents a premise with two choices to determine causality or effect; 3) the XNLI benchmark Conneau et al.
-(2018), which assesses Natural Language Inference by judging the relationship between a premise and a hypothesis; 4) the PAWS-X benchmark (Yang et al., 2019) involves identifying paraphrases in two sentences; 5) the MKQA dataset (Longpre et al., 2021), which focuses on open-domain questions; 6) the XL-Sum benchmark Hasan et al.
-(2021), which involves summarizing news articles.
-We choose a subset of 9 languages for MKQA and 5 languages for XL-Sum.
-For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks.
-More details of the benchmarks can be found in A.1.
-For each task, we sample up to 500 examples from the test set for each language.
-For MGSM, since there are only 250 examples for each language, the entire test set will be used.
-For generation tasks like MKQA and XL-Sum, we will translate the answer back to the original language if the prompting strategy involves a translator.
+For each task, we sample 500 examples from the test set per language or use the entire test set if there are fewer than 500 examples.
+For generation tasks like MKQA and XL-Sum, answers will be translated back to the original language if the prompting strategy uses a translator.
+Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
+high low high low high low high low high low high low high low
+ChatGPT Native-Basic 44.4 19.4 84.6 69.7 56.9 48.6 51.6 40.6 35.1 36.4 32.5 29.9 50.8 40.8
+EN-Basic 50.3 27.3 88.3 73.3 64.6 61.8 64.3 50.4 37.4 33.3 33.3 30.0 56.4 46.0
+Native-CoT 65.1 27.1 84.1 69.8 54.9 47.4 51.6 43.4 35.5 35.1 31.9 27.9 53.8 41.8
+EN-CoT 70.5 47.1 89.9 75.9 60.2 53.6 63.7 51.2 43.3 41.2 30.0 28.6 59.6 49.6
+XLT 70.4 50.1 89.3 76.8 60.6 58.1 59.7 58.2 37.7 37.5 22.8 26.1 56.7 51.1
+Trans-Google 74.7 72.7 90.3 83.2 62.4 59.1 68.2 62.0 42.5 48.3 30.6 28.9 61.4 59.0
+Trans-NLLB 65.6 54.1 85.7 78.2 60.5 58.2 68.4 63.4 35.4 43.6 28.4 27.7 57.3 54.2
+Llama-2-70B-Chat Native-Basic 35.7 5.6 64.2 48.0 43.0 36.0 53.3 50.4 28.9 10.4 30.1 26.8 42.5 29.5
+EN-Basic 42.5 7.7 70.7 52.0 52.7 41.9 61.9 52.8 25.7 21.5 30.2 35.3 47.3 35.2
+Native-CoT 35.5 5.6 65.3 46.8 41.0 35.6 56.0 49.6 25.3 9.9 26.0 25.2 41.5 28.8
+EN-CoT 45.6 7.0 80.7 56.3 52.7 40.9 66.5 57.0 32.7 25.7 29.8 32.0 51.3 36.5
+XLT 49.0 8.4 76.4 54.7 57.3 48.4 56.6 51.6 26.5 26.7 19.3 11.5 47.5 33.6
+Trans-Google 55.5 50.0 86.3 79.7 55.3 53.0 69.4 64.2 38.7 43.1 33.1 36.7 56.4 54.4
+Trans-NLLB 46.5 39.7 83.3 75.6 53.7 51.0 70.5 62.4 17.8 24.7 32.4 36.2 50.7 48.3
+Table 1: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting. The best result for each model is in bold.
 2.1.2 Models
-We conduct experiments with six LLMs, consisting of one closed-source language model and five open-source language models:
-  • ChatGPT [^3] https://platform.openai.com/docs/models/gpt-3-5, which is the most capable and cost-effective model in the GPT-3.5 family and it has been optimized for chat.
+We mainly conduct experiments on the following two LLMs, consisting of one closed-source language model and one open-source language model:
+ChatGPT
+This is the most capable and cost-effective model in the GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat.
 We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
-  • Mistral-7B-Instruct (v0.2). This model is the instructed version of Mistral-7B Jiang et al. (2023).
-  • Llama-2-13B-chat and Llama-2-70B-chat, which are the chat models in Llama-2 family Touvron et al.
+Llama-2-70B-Chat
+This is the largest chat models in Llama-2 family Touvron et al. (2023). Due to computational resource limitations, we use the AWQ Lin et al. (2023) version for evaluation.
+We also conducted experiments on some other models, including Mistral-7B-Instruct (v0.2) Jiang et al.
+(2023), Llama-2-13B-chat Touvron et al.
+(2023) and bloomz-7b1 Muennighoff et al.
 (2023).
-Due to computational resource limitations, we use the AWQ Lin et al.
-(2023) version for Llama-2-70B-chat.
-  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including multilingual tasks Muennighoff et al. (2023).
+More details are shown in Appendix A.1.
 2.1.3 Prompting Strategies
-Considering the language of instruction, whether to apply a chain-of-thought reasoning approach and the availability of translation tools, we assess the following experimental strategies in the experiment.
-We utilize a zero-shot approach across all strategies, given that the models we selected have been fine-tuned to follow instructions.
+Considering the language of instruction, the use of chain-of-thought reasoning, and translation tools, we assess the following experimental strategies.
+We employ a zero-shot approach for all strategies, given that our selected models are fine-tuned to follow instructions.
 Basic prompt with native instructions (Native-Basic)
@@
 Cross-lingual-thought (XLT)
-XLT Huang et al.
-(2023) is a state-of-the-art prompting method to handle multilingual NLP tasks.
-It prompts LLMs to translate the question into English and solve the problem step-by-step in English, leveraging the English-centric bias of LLMs.
+XLT Huang et al. (2023) is a state-of-the-art prompting method to handle multilingual NLP tasks. It prompts LLMs to translate the question into English and solve the problem step-by-step in English.
 Translate to English with Google Translate (Trans-Google)
@@
 Instead of using commercial translators, we use an open-source model, namely NLLB Team et al. (2022). Specifically, we chose nllb-200-3.3B to do the translation.
+The examples for each strategy are shown in Table A.1.4 and the templates for EN-Basic are shown in Table 5 in the Appendix.
 In addition to the prompting strategies, an output constraint is also included in the template to facilitate answer extraction.
-The settings are demonstrated in Table 1.
-In the case of open-source models, where the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
-2.2 Results and Discussions
-Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
-ChatGPT Native-Basic 34.4 72.3 52.2 49.7 35.4 31.9 46.0
-EN-Basic 41.1 76.1 63.0 62.0 36.5 32.7 51.9
-Native-CoT 49.9 72.4 50.6 50.3 35.4 31.1 48.3
-EN-CoT 61.1 78.6 56.4 61.6 42.9 29.7 55.1
-XLT 62.3 79.3 59.2 59.4 37.6 23.4 53.5
-Trans-Google 73.9 84.5 60.5 67.2 43.8 30.3 60.0
-Trans-NLLB 61.0 79.7 59.2 67.5 37.2 28.3 55.5
-bloomz-7b1 Native-Basic 1.3 21.4 8.3 - 7.8 - 9.7
-EN-Basic 2.0 56.5 43.9 - 10.6 - 28.3
-Native-CoT 1.2 20.9 8.2 - 6.5 - 9.2
-EN-CoT 1.7 53.9 35.9 - 9.3 - 25.2
-XLT 1.7 50.5 35.4 - 8.0 - 23.9
-Trans-Google 2.7 63.7 44.3 - 17.2 - 32.0
-Trans-NLLB 2.4 61.8 43.8 - 14.7 - 30.7
-Mistral-7B-Instruct Native-Basic 11.2 54.2 42.8 44.6 7.8 25.9 31.1
-EN-Basic 23.8 34.9 50.2 46.9 7.0 21.6 30.7
-Native-CoT 17.0 53.8 43.4 44.3 7.8 24.6 31.8
-EN-CoT 27.6 40.8 50.0 46.6 11.5 18.7 32.5
-XLT 31.8 61.5 46.0 47.8 9.6 17.6 35.7
-Trans-Google 41.3 59.2 55.0 51.5 17.0 32.8 42.8
-Trans-NLLB 31.7 54.4 53.0 52.4 15.5 32.2 39.9
-Llama-2-13b-Chat Native-Basic 15.6 50.1 36.4 54.0 18.3 27.5 33.6
-EN-Basic 19.0 54.3 43.4 59.1 20.2 31.1 37.9
-Native-CoT 18.1 50.9 35.7 54.8 15.7 27.2 33.7
-EN-CoT 19.9 54.5 43.7 57.6 19.8 30.2 37.6
-XLT 22.3 56.0 51.4 55.7 19.0 21.2 37.6
-Trans-Google 39.1 71.9 46.1 58.4 33.8 35.7 47.5
-Trans-NLLB 31.8 68.2 45.4 57.8 28.4 34.2 44.3
-Llama-2-70B-Chat Native-Basic 23.6 51.5 39.0 52.8 24.8 29.4 36.2
-EN-Basic 28.6 55.3 46.5 60.4 24.7 31.2 40.1
-Native-CoT 23.5 50.5 37.9 54.9 21.9 25.8 34.5
-EN-CoT 30.2 61.2 45.9 64.9 31.1 30.2 42.3
-XLT 32.8 58.7 52.2 55.7 26.6 17.7 38.4
-Trans-Google 53.3 80.9 54.0 68.5 39.7 33.8 51.8
-Trans-NLLB 43.8 77.1 52.2 69.2 19.4 33.1 49.1
-Table 2: Average scores across different languages (excluding English) for the six benchmarks in zero-shot setting.
-The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks.
-The best result for each model is in bold.
+When the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
+2.2 Main Results
 The main results are shown in Table 1.
-We also calculate the average performance for high-resource and low-resource languages separately for each task in Table 4 in Appendix A.2.
-We notice that Trans-Google, despite simple, surprisingly demonstrates the highest overall performance across various models and tasks.
+We notice that Trans-Google, despite simple, demonstrates the highest overall performance across various models and tasks.
 While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages.
-Besides this, we can have the following observations: 1) Trans-NLLB ranks second on average, trailing behind Trans-Google.
-The discrepancy between the two methods lies solely in the quality of translation, indicating that the NLLB model’s translation proficiency is inferior.
-Nonetheless, its open-source nature renders it a potentially favourable choice under specific circumstances.
-2) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought.
+Besides this, we can have the following observations: 1) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought.
 This finding aligns with those reported by Lai et al., 2023.
-3) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM.
+2) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM.
 For weaker models and tasks that can be answered directly, the basic prompt may be a better option.
-More results and discussions are in Appendix A.2.
+3) On average, EN-CoT underperforms compared to Trans-Google for both high and low-resource languages.
+While EN-CoT surpasses Trans-NLLB in high-resource languages, it falls short in low-resource ones.
+We hypothesize that this discrepancy arises because LLMs excel in high-resource languages but need external translation systems to handle low-resource languages effectively.
+More results and discussions are in Appendix A.1.4.
+2.3 Analysis and Discussions
+To investigate the impact of different factors on performance across various languages, we conduct a series of experiments and analyses using the MGSM benchmark.
+Is there a relationship between task performance and translation quality?
+In addition to external translation systems, we can use LLMs to translate the questions.
+Although XLT includes translation, it is integrated into the solutions.
+Therefore, we examine the self-translate approach Etxaniz et al.
+(2023), translating in a zero-shot manner with the prompt template shown in Appendix A.1.3.
+Then we prompt LLMs with the translated question the same as Trans-Google and Trans-NLLB.
+The results are shown in Table 8 in the Appendix.
+We use the English subset of MGSM as the reference translation and evaluate translation quality using the SacreBLEU score Papineni et al.
+(2002); Post (2018).
+The results, shown in Figure 3, indicate that Google Translate achieves the highest quality for all languages except Japanese.
+Translations by ChatGPT (Trans-ChatGPT) and Llama-2-70B-Chat (Trans-Llama) outperform Trans-NLLB for high-resource languages but not for some low-resource languages.
+Figure 3: BLEU scores for translating MGSM questions with different translation systems.
+To analyze the impact of translation quality on final performance, we plot the correlation between accuracy scores and BLEU scores for each language in Figure 4.
+The results show that higher translation quality (BLEU scores) generally leads to better task performance, highlighting the importance of an effective translation system.
+Figure 4: Corrections between BLEU scores of translation and MGSM accuracy for the three prompting techniques: Trans-Google, Trans-NLLB and self-translate.
+Each dot in the figure represents the performance of one model on one language.
+Does language distance between English and target language affect the performances?
+Table 1 shows that the LLMs perform better for high-resource languages than low-resource languages on average.
+We hypothesize that language distance, besides language frequency, is crucial for English-centric LLMs.
+To verify this, we calculate the correlation between MGSM accuracy and the language distances between the target languages and English.
+Following Philippy et al., 2023, We examine five types of distances, including the syntactic (SYN), geographic (GEO), inventory (INV), genetic (GEN), and phonological (PHON) distances extracted using lang2vec Littell et al.
+(2017).
+As shown in Table 2, MGSM accuracy significantly correlates with syntactic distance but not with other types of distances.
+The negative values indicate that languages with a larger syntactic distance from English tend to perform worse.
+Prompt type SYN GEO INV GEN PHON
+ChatGPT
+Native-Basic -0.786* -0.336 0.323 -0.403 -0.044
+EN-Basic -0.820* -0.160 0.527 -0.299 0.020
+Native-CoT -0.795* -0.184 0.479 -0.313 0.045
+EN-CoT -0.841* -0.286 0.339 -0.436 -0.034
+XLT -0.787* -0.113 0.445 -0.284 0.117
+Llama-2-70B-Chat
+Native-Basic -0.688* -0.369 0.250 -0.323 -0.044
+EN-Basic -0.782* -0.512 0.134 -0.513 -0.226
+Native-CoT -0.706* -0.403 0.231 -0.475 -0.105
+EN-CoT -0.737* -0.510 0.206 -0.445 -0.219
+XLT -0.697* -0.432 0.266 -0.423 -0.153
+Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language. (*p < 0.05, two-tailed)
 3 Translation for Real User Queries
+(a) Win rate with ChatGPT
+(b) Win rate with Llama-2-70B-Chat
+Figure 5: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat.
 The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages.
@@
 This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages.
-To assess the impact of translation on real-world queries, we extract user requests from ShareGPT datasets.
-We select 10 languages ranging from high to low resources and randomly sample 50 requests for each language from the first round of the dataset.
-To evaluate the quality of the responses, we use GPT-4-Turbo (gpt-4-turbo-0125) as the judge.
+To assess the impact of translation on real-world queries, we extract user requests from ShareGPT [^3] https://sharegpt.com/, a website to share real conversations with ChatGPT.
+3.1 Experiment Setup
+We selected 10 languages, ranging from high to low resource, and randomly sampled 100 requests for each language.
+However, for Romanian (ro), Ukrainian (uk), and Norwegian (no), we sampled 53, 98, and 53 requests respectively, due to the limited number of samples available from the source dataset.
+Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API.
+For the second option, we translate the output back to the original language for consistency.
+To evaluate the quality of the responses, we use GPT-4o [^4] https://openai.com/index/hello-gpt-4o/(gpt-4o-2024-05-13) as the judge.
 The prompt for the judge is shown in Figure 8 in the Appendix, which is adapted from Zheng et al.
@@
 With this prompt, each response will get a score from 1 to 10.
-Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API.
-For the second option, we will translate the output back to the original language for consistency.
-We select the two best models on NLP tasks for evaluation: ChatGPT and Llama-2-70B-Chat.
-3.1 Results and Discussions
-In our comparison, we analyzed the scores of two sets of responses generated by the same model, calculating the win rate for each language.
-The results are shown in Figure 3 and we make the following observations: 1) ChatGPT shows different behaviours for high-resource languages and low-resource languages.
-For Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate.
-For lower-resource languages, translation is often a better option.
-2) For Llama-2-70B-Chat, translation has a higher win rate for all the languages.
-This echoes the fact that Llama-2-70B-Chat is an English-centric LLM.
-Even though translation may lose some information, the gain of better understanding after translation dominates the performance.
-The distinct behaviours exhibited by Llama-2-70B-Chat and ChatGPT underscore the inherent differences between the LLMs.
-Llama-2-70B-Chat is predominantly English-centric, as evidenced by its improved performance in both NLP tasks and real user queries when input is translated into English.
-Conversely, ChatGPT demonstrates characteristics of a “multilingual LLM”, as shown in Figure 1(b), showcasing its broader linguistic capacity.
-Nevertheless, ChatGPT still relies on translation to bolster its performance for lower-resource languages, indicating room for improvement in its approach to truly multilingual processing.
-To ascertain the extent to which answering user queries necessitates an understanding of local cultural nuances, we employed GPT-4-Turbo, utilizing a specially crafted prompt to analyze user queries across multiple languages, as shown in Figure 7 in the Appendix.
-The results, illustrated in Figure 4, reveal that a significant portion of the queries—ranging from 18% to 50% for each language—require knowledge of local culture for accurate responses.
-This variation underscores the rich cultural elements embedded within the data, affirming our hypothesis that queries containing culture-specific information pose a greater challenge for the LLMs.
-(a) Win rate with ChatGPT
-(b) Win rate with Llama-2-70B-Chat
-Figure 3: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat.
-3.2 Analysis
-Figure 4: The percentage of the questions that necessitate local cultural knowledge.
-ChatGPT has shown promising performance in some high-resource languages.
-We extend the experiment to investigate the effectiveness of multilingual LLMs in handling culture-specific tasks, aiming to assess how well these models grasp cultural nuances and idiomatic language.
-Through our experiment, we seek insights into the adaptability of LLMs across diverse cultural contexts.
-3.2.1 Experiment Setup
-We select the following models to represent various categories of LLMs and assess the potential differences in their response to translation:
-  • ChatGPT and llama-2-70B-chat, which are the strongest models in our previous experiments.
-  • Qwen1.5-72B-Chat Bai et al.
-(2023) and Yi-34B-Chat [^4] https://github.com/01-ai/Yi.
-These two open-source models demonstrate strong English and Chinese capabilities.
-We aim to test if they excel more in Chinese than in translating Chinese into English.
-Please note that except ChatGPT, the results for the other three models are obtained by calling API from Together [^5] https://api.together.xyz/.
-We chose open-source models which are larger than 30B since we hypothesise that larger models contain more knowledge.
-For the evaluation dataset, we choose M3Exam Zhang et al.
-(2023a), as the questions require strong multilingual proficiency and cultural knowledge to perform well.
-For the subject categories, we choose language and social science, which are more likely to contain language-dependent knowledge.
-For each language, we select up to 500 samples for evaluation.
-For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
-3.2.2 Results and Discussions
 Figure 6: Results of four LLMs on M3Exam (a) language and (b) social science subject categories.
 Note that in M3Exam, not all subjects are available in every language, resulting in a discrepancy in the number of languages for the two subjects.
-Based on the results shown in Figure 5, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance for all languages.
+3.2 Main Results
+We compared the scores of two response sets from the same model, calculating the win rate for each language.
+The results are shown in Figure 5, leading to the following observations: 1) ChatGPT’s performance varies across languages.
+For high-resource languages like Japanese, Chinese, and Spanish, original queries have a higher win rate.
+In contrast, for low-resource languages, the effectiveness of translation can be either better or worse, depending on the specific languages involved.
+2) For Llama-2-70B-Chat, translation has a higher win rate for all languages, reflecting its English-centric nature.
+Despite potential information loss, the improved understanding after translation still enhances performance.
+Llama-2-70B-Chat and ChatGPT exhibit distinct behaviors, reflecting their inherent differences.
+Llama-2-70B-Chat, being English-centric, performs better with translated inputs.
+Conversely, ChatGPT shows certain characteristics of a “Multilingual LLM”, as shown in Figure 1(b), mainly for high-resource languages, indicating the potential for improvement in true multilingual processing.
+To determine if answering user queries requires local cultural knowledge, we used GPT-4o with a specially crafted prompt to analyze queries in multiple languages (Figure 9 in the Appendix).
+Results in Table 14 in the Appendix show that 30% to 74% of queries per language require cultural knowledge, highlighting the rich cultural elements in the data.
+Further analysis of the ShareGPT subsets requiring local cultural knowledge is in Appendix A.2.
+3.3 Analysis and Discussions
+Based on the previous results, ChatGPT and Llama-2-70B-chat both tend to be English-centric but ChatGPT demonstrates certain behaviors of a "Multilingual LLM".
+An ideal "Multilingual LLM" should exhibit equivalent performance across all languages.
+Consequently, we broaden our analysis to include non-English-centric LLMs and assess their performance across various tasks.
+How do non-English-centric LLMs perform on culture-related tasks?
+To investigate the behaviors of different LLMs on culture-related tasks, we select another two LLMs: Qwen1.5-72B-Chat Bai et al.
+(2023) and Yi-34B-Chat AI et al.
+(2024), which are not English-centric.
+These two open-source models demonstrate strong capabilities in both English and Chinese.
+Therefore, we can check whether they demonstrate multilingual behaviors in Chinese, as illustrated in Figure 1(b).
+For the evaluation dataset, we choose M3Exam Zhang et al.
+(2023a), as the questions are real-world natural data from different languages instead of translating from English data and require strong multilingual proficiency and cultural knowledge to perform well.
+For example, the question about a Swahili proverb shown in Figure 2(b) requires local knowledge to answer correctly.
+We select the language and social science subject categories, which likely contain more native cultural knowledge, and evaluate up to 500 samples per language.
+For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
+Based on the results shown in Figure 6, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance.
 This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a.
@@
 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency.
-The translation process hurts Chinese performance, highlighting the significant influence of translation errors on comprehension.
-Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat.
-From this experiment, it is evident that while translation serves as a straightforward and initially effective strategy to enhance the performance of English-centric LLMs on non-English tasks, it falls short of being optimal.
-Directly training LLMs in specific languages can yield superior results by eliminating the potential for translation errors.
-4 General Discussions
-In this work, we categorize multilingual tasks into two groups: language-independent tasks and language-dependent tasks.
-Among the NLP benchmarks discussed in Section 2, such as MGSM, XCOPA, XNLI, PAWS-X, and MKQA, are parallel benchmarks.
-This means that the examples maintain semantic equivalence across all languages, categorizing them as language-independent tasks.
-Language-independent tasks are typically easier to evaluate across different linguistic groups, as they rely on universal cognitive abilities, like reasoning and comprehension, rather than specific linguistic knowledge.
-In contrast, tasks involving real user queries and culture-related questions, discussed in Section 3, necessitate models to possess cultural and contextual understanding, categorizing them as language-dependent tasks.
-By distinguishing between language-dependent and language-independent tasks, we aim to promote fair and accurate evaluations across diverse linguistic contexts, enhancing the development and assessment of NLP systems in multilingual environments.
-Corresponding to the two types of tasks, multilingual LLMs must develop dual competencies:
-General Language Skills:
-Essential for understanding and engaging with both task types, these skills enable LLMs to process and comprehend content across multiple languages, facilitating basic instruction interpretation and response generation.
-Specific Language Skills:
-These skills are crucial for tasks demanding an understanding of cultural nuances and linguistic subtleties, such as interpreting culturally specific queries.
-By equipping LLMs with these comprehensive skills, we ensure their effectiveness in a wide array of multilingual contexts, enhancing both linguistic adaptability and cultural sensitivity.
-The benefits of translation on NLP tasks demonstrate the dominance of English-centric LLMs.
-Even top models like ChatGPT, perform better through translation.
-In handling user queries, ChatGPT excels in native languages compared to translation, particularly in high-resource languages but struggles with low-resource ones.
-When testing on the M3Exam questions, Qwen1.5-72B-Chat and Yi-34B-Chat perform quite well in the native language for Chinese questions but have no such advantages for other languages.
-This dependency on translation highlights the challenges in developing LLMs that can equally handle a wide range of languages, especially those with limited resources.
-Such findings prompt further research and development efforts aimed at enhancing the multilingual efficacy of LLMs, striving for models that can understand and generate text across languages with minimal reliance on translation.
+The translation process hurts Chinese performance, highlighting the significant influence of translationese on comprehension.
+Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
+How do non-English-centric LLMs perform on NLP tasks?
+Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
+Qwen1.5-72B-Chat
+Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9
+EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4
+Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1
+EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7
+XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9
+Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2
+Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5
+Yi-34B-Chat
+Native-Basic 63.2 92.6 46.0 43.6 13.4 36.9 49.3
+EN-Basic 66.8 93.6 52.6 74.6 15.5 35.1 56.4
+Native-CoT 65.2 91.8 42.6 43.6 13.0 36.6 48.8
+EN-CoT 70.0 93.6 48.2 74.8 12.1 33.1 55.3
+XLT 56.0 93.2 69.2 65.6 7.5 31.3 53.8
+Trans-Google 71.2 94.0 49.6 70.8 24.5 36.3 57.7
+Trans-NLLB 56.0 86.6 48.8 68.2 22.9 28.5 51.8
+Table 3: Scores of the two non-English-centric LLMs on NLP tasks for the Chinese language. The best result for each model is in bold.
+As shown in Figure 2(b), for an ideal multilingual LLM, prompting in native languages should still have advantages over translation if the tasks are less dependent on languages.
+To test the hypothesis, we evaluate Qwen1.5-72B-Chat and Yi-34B-Chat on the NLP tasks as discussed in Section 2.1.1.
+We only evaluate them in Chinese since the two models are optimized for this language.
+The results are displayed in Table 3.
+Trans-Google remains competitive among various prompting strategies, achieving the best average scores for Yi-34B-Chat, which surpasses our expectations.
+The possible reason could be that while both models are optimized for Chinese, their performance in Chinese still lags behind their proficiency in English.
+Nevertheless, We have the following special observations for the two models.
+1) For Qwen1.5-72B-Chat, the best strategy is EN-CoT instead of Trans-Google.
+We hypothesize that this prompting strategy utilizes the model’s bilingual abilities and simultaneously avoids translationese.
+2) Both LLMs perform better with Native-Basic for the XL-Sum dataset.
+We hypothesize that the dataset is more language-dependent than other tasks as it is created by considering the local context instead of simply translating from the English version Hasan et al.
+(2021).
+3) The translation benefits are less pronounced than those of ChatGPT and Llama-2-70B-Chat.
+For example, the gap between Trans-Google and Native-Basic on MGSM(Chinese) for the two models are 2.8% and 8%.
+The values for ChatGPT and Llama-2-70b-Chat are 37.2% and 16%, respectively, which are significantly larger.
+How do different LLMs handle multilingual prompts?
+To further understand the differences between English-centric LLMs and non-English-centric LLMs, we analyze the layerwise language distribution for Llama-2-7B-Chat and Qwen1.5-7B-Chat, using the method proposed by Zhao et al., 2024.
+We decode the embedding after each layer and identify each token into different languages with CLD3555https://github.com/google/cld3.
+As shown in Figure 7, the two LLMs process Chinese prompts differently.
+While the hidden representations of Qwen1.5-7B-Chat are mainly in Chinese, those of Llama-2-7B-Chat are in various other languages.
+We hypothesize that processing the information in native without conversion avoids the information loss, making it more suitable for processing multilingual tasks.
+(a) Llama-2-7B-chat
+(b) Qwen1.5-7B-Chat
+Figure 7: Layerwise language distribution for (a) Llama-2-7b-chat and (b) Qwen1.5-7B-Chat with Chinese prompts.
+4 Related Work
+Multilingual Evaluation.
+Since the release of ChatGPT, the evaluation of LLMs has attracted the attention of the research communityQin et al.
+(2023a); Bang et al.
+(2023).
+Shi et al., 2022 evaluated LLMs on MGSM and found that the models demonstrated strong multilingual reasoning capabilities, even for low-resource languages.
+Bang et al., 2023 evaluated ChatGPT on 23 datasets covering 8 NLP tasks.
+They found that ChatGPT failed to generalize its capabilities to non-Latin scripts.
+To cover tasks, Ahuja et al., 2023 evaluated ChatGPT and GPT-4 on 16 NLP datasets across 70 languages and compared them with state-of-the-art non-autoregressive models.
+Concurrently, Lai et al., 2023 evaluated ChatGPT on 7 different tasks across 37 diverse languages.
+However, these evaluations are primarily limited to standard NLP tasks and largely overlook real-world scenarios and cultural knowledge Fung et al.
+(2024), which are crucial for understanding the practical applicability of LLMs.
+Multilingual Prompting Strategies.
+The translate-test is a popular technique used to refine the performance of multilingual NLP benchmarks Conneau et al.
+(2018); Ponti et al.
+(2020); Artetxe et al.
+(2023); Moghe et al.
+(2023).
+In the era of LLMs, various strategies have been developed to enhance the performance of LLMs using multilingual datasets.
+Shi et al., 2022 discovered that EN-CoT outperforms Native-CoT.
+Huang et al., 2023 introduced cross-lingual-thought prompting to minimize language disparities.
+In parallel, Qin et al., 2023b introduced cross-lingual prompting, and Etxaniz et al., 2023 suggested self-translate to elevate their performances.
+Effective in translating prompts into English, these methods excel in NLP tasks but remain uncertain in real-world applications.
+Their success hinges on the English-centric nature of the LLMs.
+Our study evaluates translation effectiveness across NLP tasks, real user queries, and non-English-centric LLMs, revealing the limitations of these methods.
 5 Conclusion
-In this work, we perform a comprehensive assessment of LLMs across diverse multilingual tasks.
-These tasks include traditional NLP challenges, real user queries, and culture-specific tasks.
-We explore the use of translation-based methods as a strategy to overcome the limitations inherent in English-centric LLMs, thereby facilitating enhanced comprehension.
-However, we find that for tasks heavily reliant on language-specific nuances, employing prompts in the native language emerges as a more effective approach.
+We have conducted a thorough evaluation of LLMs in various multilingual tasks.
+These tasks include traditional NLP benchmarks, real user queries, and culture-related tasks.
+Even though translation-based methods are simple and effective strategies to overcome the limitations inherent in English-centric LLMs, they are not optimal for all scenarios, highlighting the necessity of more comprehensive multilingual evaluation.
+The experiment on non-English-centric LLMs and culture-related tasks demonstrates that employing prompts in the native language emerges as a more effective approach.
 This method is particularly adept at capturing the subtleties and intricacies unique to each language.
-Our findings suggest that the development of robust multilingual LLMs presents a more effective solution for addressing the challenges of multilingual tasks, especially for language-dependent tasks.
-Therefore, we advocate for the prioritization of research and development efforts towards the creation of strong multilingual large language models.
-6 Limitations
+The challenge of the setting is that it requires LLMs to be proficient in various languages, calling for the prioritization of research and development efforts toward the creation of strong multilingual LLMs.
+Limitations
 This study aims to systematically assess the effectiveness of various prompting strategies across different tasks and LLMs.
@@
 This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore.
-Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and School of Computer Science and Engineering, Nanyang Technological University, Singapore, for their support.
-Appendix A Additional Experiments
-A.1 NLP Tasks
-Here are the detailed descriptions of the NLP tasks:
+Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and College of Computing and Data Science, Nanyang Technological University, Singapore, for their support.
+Appendix A Appendix
+A.1 Translation for NLP Tasks
+This section presents more details about the setups and results for the experiments on NLP tasks.
+A.1.1 Details about NLP Benchmarks
+Here are the detailed descriptions of the NLP benchmarks:
 Arithmetic Reasoning
@@
 We use the ROUGE-1 score for evaluation.
+A.1.2 More LLMs for Experiment
+Besides ChatGPT and Llama-2-70B-Chat, we have also evaluated the NLP tasks with the following models:
+  • Mistral-7B-Instruct (v0.2). This model is the instructed version of Mistral-7B Jiang et al. (2023).
+  • Llama-2-13B-chat, which is a chat model in Llama-2 family Touvron et al. (2023).
+  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including some multilingual tasks Muennighoff et al. (2023).
+A.1.3 More Details about Prompt Strategies
+An example of various prompting strategies is shown in Table A.1.4.
 The prompts of EN-Basic for each task are shown in Table 5, which are adapted from Huang et al., 2023.
+The translation template for self-translate with LLMs is:
+Translate the following question from {language} to English:
+{question}
+Don’t answer the question, just translate it!
 The prompt templates for other prompting strategies and the instructions for output formats are designed according to the descriptions in Section 2.1.3.
-A.2 Additional Results
+A.1.4 Additional Results
 The average performances for high-resource and low-resource languages are shown in Table 6.
@@
 For open-source models, while we did not observe a refusal pattern, they do not follow the instructions properly, which also degrades their performance with XLT.
+Original Question {CJK*}UTF8gbsn制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
+Native-Basic {Original Question }
+{CJK*}UTF8gbsn您的最终答案的格式应为:"答案: <阿拉伯数字>".
+EN-Basic {Original Question }
+You should format your final answer as "Answer: <Arabic numeral>".
+Native-CoT {Original Question }
+{CJK*}UTF8gbsn让我们一步步思考。
+{CJK*}UTF8gbsn您的最终答案的格式应为:"答案: <阿拉伯数字>".
+EN-CoT {Original Question }
+Let’s think step by step in English.
+You should format your final answer as "Answer: <Arabic numeral>".
+XLT I want you to act as an arithmetic reasoning expert for Chinese.
+Request: {Original Question}
+You should retell the request in English.
+You should do step-by-step answer to obtain a number answer.
+You should step-by-step answer the request.
+You should tell me the answer in this format ’Answer :’.
+Trans-Google Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
+Let’s think step by step.
+You should format your final answer as "Answer: <Arabic numeral>".
+Trans-NLLB To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
+Let’s think step by step.
+You should format your final answer as "Answer: <Arabic numeral>".
+Table 4: An example of zero-shot prompts for a Chinese problem.
+For Native-Basic, EN-Basic, Native-CoT, EN-CoT and XLT, we provide the original Chinese question as input and expect an answer in the corresponding format; for Trans-Google and Trans-NLLB, we input the translated question in English, and expect a step-by-step solution in English.
+To obtain the desirable output format, we instruct the models to output in specific format.
 Benchmark #Test Basic Prompt
 MGSM 250 {problem}
-XCOPA 500
-Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
-XNLI 500
-{premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
-PAWS-X 500
-Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
-MKQA 500
-Answer the question in one or a few words in {target_language}: {question}?
-XL-Sum 500
-{article} Summarize the article.
+XCOPA 500 Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
+XNLI 500 {premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
+PAWS-X 500 Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
+MKQA 500 Answer the question in one or a few words in {target_language}: {question}?
+XL-Sum 500 {article} Summarize the article.
 Table 5: Template of EN-Basic for each benchmark. #Test denotes the number of samples in the test set.
@@
 Table 6: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting.
+The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks.
+The best result for each model is in bold.
 Model Prompt type de ru fr zh es ja sw th bn te avg
@@
 Table 7: Accuracy scores across various languages on the MGSM benchmark.
+Model Prompt type de ru fr zh es ja sw th bn te avg
+ChatGPT Trans-ChatGPT 77.6 75.2 78.4 76.0 78.8 69.6 75.2 62.4 65.6 42.8 70.2
+Llama-2-70B-Chat Trans-Llama 53.6 52.0 55.2 46.8 54.4 44.0 8.8 11.2 15.2 4.8 34.6
+Table 8: Accuracy scores across various languages on the MGSM benchmark with self-translate approach.
 Model Prompt type zh it vi tr id sw th et ta ht qu avg
@@
 Table 13: ROUGE-1 scores across various languages on the XL-sum benchmark.
-Figure 6: The LLM-as-a-judge prompt for GPT-4-Turbo.
-Figure 7: Prompt to check whether answering a request needs local cultural knowledge.
+A.2 Translation for Real User Queries
+The prompt used to assess the response quality is shown in Figure 8.
+When GPT-4o is prompted with this, it assigns a score ranging from 1 to 10 to each response.
+Figure 9 illustrates the prompt used to determine if responding to a request requires local cultural knowledge.
+The Chinese case shows that GPT-4o can identify if queries require knowledge of local culture with explanations and the final answer.
+We also analyzed the performance of shareGPT subsets with cultural knowledge only.
+As shown in Figure 10, the behaviors across languages and models are inconsistent.
+ChatGPT shows different behaviors for high-resource and low-resource languages.
+For high-resource languages like Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate.
+For low-resource languages, translation is often a better option.
+In contrast, Llama-2-70B-Chat shows a higher win rate for all languages.
+Figure 8: The LLM-as-a-judge prompt for GPT-4o.
+Figure 9: Prompt template to check whether answering a request needs local cultural knowledge (upper) and one Chinese example (lower).
+Language ja zh es fr vi id ko ro uk no
+Ratio (%) 59 58 38 41 67 55 55 74 30 57
+Table 14: The percentage of the questions that necessitate local cultural knowledge.
+(a) Win rate with ChatGPT (w/ cultural knowledge)
+(b) Win rate with Llama-2-70B-Chat (w/ cultural knowledge)
+Figure 10: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat for the subsets of shareGPT with cultural knowledge.

### v3

@@
 Abstract
-Large language models (LLMs) have demonstrated multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora.
-Existing works leverage this phenomenon to improve their multilingual performances through translation, primarily on natural language processing (NLP) tasks.
-This work extends the evaluation from NLP tasks to real user queries and from English-centric LLMs to non-English-centric LLMs.
-While translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios.
-For culture-related tasks that need deep language understanding, prompting in the native language tends to be more promising as it better captures the nuances of culture and language.
-Our experiments reveal varied behaviors among different LLMs and tasks in the multilingual context.
-Therefore, we advocate for more comprehensive multilingual evaluation and more efforts toward developing multilingual LLMs beyond English-centric ones.
+Large language models (LLMs) have demonstrated multilingual capabilities, yet they are mostly English-centric due to the imbalanced training corpora.
+While prior works have leveraged this bias to enhance multilingual performance through translation, they have been largely limited to natural language processing (NLP) tasks.
+In this work, we extend the evaluation to real-world user queries and non-English-centric LLMs, offering a broader examination of multilingual performance.
+Our key contribution lies in demonstrating that while translation into English can boost the performance of English-centric LLMs on NLP tasks, it is not universally optimal.
+For culture-related tasks that need deep language understanding, prompting in the native language proves more effective as it better captures the nuances of culture and language.
+Our experiments expose varied behaviors across LLMs and tasks in the multilingual context, underscoring the need for a more comprehensive approach to multilingual evaluation.
+Therefore, we call for greater efforts in developing and evaluating LLMs that go beyond English-centric paradigms.
+[^1] Our code is publicly available at https://github.com/DAMO-NLP-SG/translation-all-you-need.
 Is Translation All You Need?
@@
 12  Wenxuan Zhangthanks: Wenxuan Zhang is the corresponding author.
-23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bing23 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore {chaoqun.liu,saike.zwx,zhaoyiran.zyr,l.bing}@alibaba-inc.com anhtuan.luu@ntu.edu.sg
+23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bingthanks: Work done while at Alibaba Group.
+5 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore; 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore; 5Shanda AI Research Institute {chaoqun.liu,saike.zwx}@alibaba-inc.com; lidong.bing@shanda.com
 1 Introduction
@@
 (2022); Workshop et al.
+(2023); Lin et al.
 (2022), resulting in an English-centric bias in LLMs.
@@
 (2023); Moghe et al.
+(2023); Bareiß et al.
 (2024).
@@
 Despite these advancements, methodologies in various studies differ significantly, and the impact of translation on multilingual task performance remains underexplored.
-Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, with less attention given to real-world user requests in various languages.
+Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, but did not study real-world user queries in various languages.
 This gap highlights a need for more nuanced research into the effectiveness of translation techniques across multilingual contexts.
 As shown in Figure 1, we hypothesize that English-centric LLMs generally perform better with English translations of prompts, while "Multilingual LLMs" excel with native prompts, particularly for tasks highly dependent on language.
-In this study, we perform an in-depth analysis of the utility of translation with large language models for various scenarios.
+To address the limitations of existing empirical studies, we perform an in-depth analysis of the utility of translation with large language models for various scenarios.
 Firstly, we compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages (i.e., the “translate-test” method), against several baselines on multilingual NLP tasks.
@@
 Thirdly, we broaden the scope of LLM evaluations to include non-English-centric models to explore how they differ in behavior from English-centric LLMs.
+To the best of our knowledge, this is the first work to analyze the impacts of translating real user queries on multilingual LLMs.
 Our results demonstrate that simply translating queries into English can already achieve the best results in multiple NLP task categories.
@@
 In addition, the non-English-centric LLMs also behave differently from English-centric LLMs, where prompts in the native languages yield superior results by capturing the nuances related to culture and language.
-The main contributions of this work can be summarized as follows:
-  • We compare various multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs.
-Our study also identifies factors that affect performance in multilingual tasks.
-  • We broaden our multilingual evaluation to encompass actual user queries and LLMs that are not primarily centered on English, thereby addressing the limitations of previous works.
-  • Our study reveals the shortcomings of current multilingual evaluations, underscoring the need for more comprehensive assessments that broaden the range of benchmarks and diversify the types of LLMs considered.
+The main contributions of this work are:
+  • We conduct a comprehensive comparison of multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs, and identifying factors impacting multilingual performance.
+  • We expand multilingual evaluation to include actual user queries and and non-English-centric LLMs, addressing the limitations of previous studies.
+  • We expose critical gaps in current multilingual evaluations, underscoring the need for more comprehensive benchmarks and a broader range of LLMs.
 2 Translation for NLP Tasks
@@
 ChatGPT
-This is the most capable and cost-effective model in the GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat.
-We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
+This is the most capable and cost-effective model in the GPT-3.5333https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat.
+We chose the latest version (gpt-3.5-turbo-1106) for the experiment.
 Llama-2-70B-Chat
@@
 2.1.3 Prompting Strategies
-Considering the language of instruction, the use of chain-of-thought reasoning, and translation tools, we assess the following experimental strategies.
-We employ a zero-shot approach for all strategies, given that our selected models are fine-tuned to follow instructions.
+We assess experimental strategies based on language of instruction, chain-of-thought reasoning, and translation tools, using a zero-shot approach as the selected models are fine-tuned for instruction-following.
 Basic prompt with native instructions (Native-Basic)
-This is to directly pose the questions without employing additional prompting strategies such as chain-of-thought.
-Both the query information and the instruction provided are presented in the original language.
+The questions are posed directly without using prompting strategies like chain-of-thought. Both the query and instructions are presented in their original language.
 Basic prompt with English instructions (EN-Basic)
@@
 We hypothesize that this discrepancy arises because LLMs excel in high-resource languages but need external translation systems to handle low-resource languages effectively.
-More results and discussions are in Appendix A.1.4.
+These findings are also applicable to smaller models, such as Mistral-7B-Instruct, as demonstrated in Table 6 in the Appendix.
+This suggests that the observations generalize well across different model types and sizes.
+Further results and discussions are provided in Appendix A.1.4.
 2.3 Analysis and Discussions
@@
 XLT -0.697* -0.432 0.266 -0.423 -0.153
-Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language. (*p < 0.05, two-tailed)
+Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language.
+A lower value indicates higher correlation due to the negative coefficients.(*p < 0.05, two-tailed)
 3 Translation for Real User Queries
@@
 Figure 5: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat.
-The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages.
-However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension.
-Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al.
+NLP tasks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries which cover diverse topics and require nuanced comprehension.
+Moreover, these benchmarks are often constructed by translating from the English data Shi et al.
 (2022); Ponti et al.
@@
 However, for Romanian (ro), Ukrainian (uk), and Norwegian (no), we sampled 53, 98, and 53 requests respectively, due to the limited number of samples available from the source dataset.
-Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API.
+Since the queries can be in various formats, we only compare two prompting strategies: 1) original queries; and 2) translated queries with Google Translate API.
 For the second option, we translate the output back to the original language for consistency.
@@
 With this prompt, each response will get a score from 1 to 10.
-Figure 6: Results of four LLMs on M3Exam (a) language and (b) social science subject categories.
-Note that in M3Exam, not all subjects are available in every language, resulting in a discrepancy in the number of languages for the two subjects.
+Figure 6: Accuracies of four LLMs on M3Exam (a) language and (b) social science subject categories.
+In M3Exam, not all subjects are available in every language, causing a difference in language coverage between the two subjects.
 3.2 Main Results
@@
 Further analysis of the ShareGPT subsets requiring local cultural knowledge is in Appendix A.2.
+We also conduct additional experiments, detailed in Appendix A.2.1, to verify that advanced LLMs can reliably assess the quality of responses.
 3.3 Analysis and Discussions
 Based on the previous results, ChatGPT and Llama-2-70B-chat both tend to be English-centric but ChatGPT demonstrates certain behaviors of a "Multilingual LLM".
-An ideal "Multilingual LLM" should exhibit equivalent performance across all languages.
 Consequently, we broaden our analysis to include non-English-centric LLMs and assess their performance across various tasks.
@@
 For the evaluation dataset, we choose M3Exam Zhang et al.
-(2023a), as the questions are real-world natural data from different languages instead of translating from English data and require strong multilingual proficiency and cultural knowledge to perform well.
-For example, the question about a Swahili proverb shown in Figure 2(b) requires local knowledge to answer correctly.
+(2023a), as the questions are real-world natural data from different languages instead of translating from English and require strong multilingual proficiency and cultural knowledge to perform well.
+For example, the question about a Swahili proverb in Figure 2(b) requires local knowledge to answer correctly.
 We select the language and social science subject categories, which likely contain more native cultural knowledge, and evaluate up to 500 samples per language.
-For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
 Based on the results shown in Figure 6, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance.
-This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a.
+This observation aligns with the conclusions in the study by Zhang et al., 2023a.
 The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension.
@@
 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency.
-The translation process hurts Chinese performance, highlighting the significant influence of translationese on comprehension.
-Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
+The translation hurts Chinese performance, highlighting the significant influence of translationese on comprehension.
+Despite this, it may boost performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
 How do non-English-centric LLMs perform on NLP tasks?
-Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
-Qwen1.5-72B-Chat
-Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9
-EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4
-Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1
-EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7
-XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9
-Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2
-Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5
-Yi-34B-Chat
-Native-Basic 63.2 92.6 46.0 43.6 13.4 36.9 49.3
-EN-Basic 66.8 93.6 52.6 74.6 15.5 35.1 56.4
-Native-CoT 65.2 91.8 42.6 43.6 13.0 36.6 48.8
-EN-CoT 70.0 93.6 48.2 74.8 12.1 33.1 55.3
-XLT 56.0 93.2 69.2 65.6 7.5 31.3 53.8
-Trans-Google 71.2 94.0 49.6 70.8 24.5 36.3 57.7
-Trans-NLLB 56.0 86.6 48.8 68.2 22.9 28.5 51.8
+Prompt type Qwen1.5-72B-Chat Yi-34B-Chat
+MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
+Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9 63.2 92.6 46.0 43.6 13.4 36.9 49.3
+EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4 66.8 93.6 52.6 74.6 15.5 35.1 56.4
+Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1 65.2 91.8 42.6 43.6 13.0 36.6 48.8
+EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7 70.0 93.6 48.2 74.8 12.1 33.1 55.3
+XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9 56.0 93.2 69.2 65.6 7.5 31.3 53.8
+Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2 71.2 94.0 49.6 70.8 24.5 36.3 57.7
+Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5 56.0 86.6 48.8 68.2 22.9 28.5 51.8
 Table 3: Scores of the two non-English-centric LLMs on NLP tasks for the Chinese language. The best result for each model is in bold.
@@
 To further understand the differences between English-centric LLMs and non-English-centric LLMs, we analyze the layerwise language distribution for Llama-2-7B-Chat and Qwen1.5-7B-Chat, using the method proposed by Zhao et al., 2024.
-We decode the embedding after each layer and identify each token into different languages with CLD3555https://github.com/google/cld3.
+We decode the embedding after each layer and identify each token into different languages with CLD3666https://github.com/google/cld3.
 As shown in Figure 7, the two LLMs process Chinese prompts differently.
@@
 We hypothesize that processing the information in native without conversion avoids the information loss, making it more suitable for processing multilingual tasks.
+In addition, we examine the layerwise language distribution in larger models, specifically Llama-2-70B-Chat and Qwen1.5-72B-Chat, as shown in Figure 12 within Appendix A.3.
 (a) Llama-2-7B-Chat
@@
 (2023); Moghe et al.
+(2023); Qi et al.
+(2022); Huang et al.
 (2022).
@@
 Acknowledgements
-This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore.
-Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and College of Computing and Data Science, Nanyang Technological University, Singapore, for their support.
+This research is supported, in part, by DSO Singapore under the research grant DSOCL23216.
+This research is also supported by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore.
+Chaoqun Liu extends his gratitude to Interdisciplinary Graduate Programme and College of Computing and Data Science of NTU, for their support.
+We sincerely appreciate the valuable feedback from Hou Pong Chan (DAMO Academy, Alibaba Group).
 Appendix A Appendix
@@
 In contrast, Llama-2-70B-Chat shows a higher win rate for all languages.
+A.2.1 Additional Results
+In Section 3.1, we randomly select 100 requests for each language and evaluate the quality of the responses generated by GPT-4o.
+To ensure a more rigorous and comprehensive analysis, we conduct additional experiments under the following conditions: we heuristically filter queries using GPT-4o to ensure their validity, select 200 queries per language from the filtered set, and employ multiple judge models.
+Due to an insufficient number of available queries in other languages, we limit our evaluation to Japanese (ja), Chinese (zh), Spanish (es), French (fr), and Korean (ko).
+For the judging process, we use not only GPT-4o but also Claude-3.5-Sonnet and Gemini-Pro-1.5 to provide a more diverse assessment.
+The results are presented in Figure 11.
+ChatGPT performs better when given direct prompts in languages such as Japanese and Chinese, whereas Llama-2-70B-Chat consistently achieves higher performance with translated prompts.
+These findings align with those discussed in Section 3.2.
 Figure 8: The LLM-as-a-judge prompt for GPT-4o.
@@
 Figure 10: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat for the subsets of shareGPT with cultural knowledge.
+(a) ChatGPT judged by GPT-4o
+(b) Llama-2-70B-Chat judged by GPT-4o
+(c) ChatGPT judged by Claude-3.5-Sonnet
+(d) Llama-2-70B-Chat judged by Claude-3.5-Sonnet
+(e) ChatGPT judged by Gemini-1.5-Pro
+(f) Llama-2-70B-Chat judged by Gemini-1.5-Pro
+Figure 11: Win rate comparison for five languages using ChatGPT and Llama-2-70B-Chat judged with three advanced LLMs.
+A.3 Layerwise Language Distribution in Larger Model
+Figure 12 illustrates the layerwise language distribution in larger models, including Llama-2-70B-Chat and Qwen1.5-72B-Chat.
+Llama-2-70B-Chat exhibits the same phenomenon as its smaller counterpart, Llama-2-7B-chat, with diverse languages represented in its hidden states.
+In contrast to Qwen1.5-7B-Chat, the hidden representations of Qwen1.5-72B-Chat incorporate both Chinese and English until the last few layers, possibly reflecting the challenges of constructing such a large model using Chinese exclusively for hidden representations.
+Nevertheless, it still represents its hidden states more in Chinese than Llama-2-70B-Chat.
+(a) Llama-2-70B-Chat
+(b) Qwen1.5-72B-Chat
+Figure 12: Layerwise language distribution for (a) Llama-2-70b-Chat and (b) Qwen1.5-72B-Chat with Chinese prompts.